### v1

# Safety Mirage: How Spurious Correlations Undermine VLM Safety Fine-tuning

Abstract
Recent vision-language models (VLMs) have made remarkable strides in generative modeling with multimodal inputs, particularly text and images. However, their susceptibility to generating harmful content when exposed to unsafe queries raises critical safety concerns. While current alignment strategies primarily rely on supervised safety fine-tuning with curated datasets, we identify a fundamental limitation we call the “safety mirage” where supervised fine-tuning inadvertently reinforces spurious correlations between superficial textual patterns and safety responses, rather than fostering deep, intrinsic mitigation of harm. We show that these spurious correlations leave fine-tuned VLMs vulnerable even to a simple one-word modification-based attack, where substituting a single word in text queries with a spurious correlation-inducing alternative can effectively bypass safeguards. Additionally, these correlations contribute to the over prudence, causing fine-tuned VLMs to refuse benign queries unnecessarily. To address this issue, we show machine unlearning (MU) as a powerful alternative to supervised safety fine-tuning as it avoids biased feature-label mappings and directly removes harmful knowledge from VLMs while preserving their general capabilities. Extensive evaluations across safety benchmarks show that under one-word attacks, MU-based alignment reduces the attack success rate by up to 60.17% and cuts unnecessary rejections by over 84.20%. Codes are available at https://github.com/OPTML-Group/VLM-Safety-MU. WARNING: There exist AI generations that may be offensive in nature.
1 Introduction
Recent multi-modal models have achieved advancements by integrating text and images [1, 2, 3, 4, 5]. A prevalent architecture in VLMs maps visual embeddings into the language model’s latent space via a dedicated projection module [6, 7, 8, 9, 10]. While the large language model (LLM) backbone effectively integrates projected visual information, the image modality also introduces new vulnerabilities [11, 12, 13, 14, 15]. Specifically, images can function as a “foreign language” [16], creating pathways for unsafe input queries, even when the underlying LLM has been aligned for safety [16, 17, 18]. The above highlights the unique safety alignment challenges that VLMs face, distinguishing them from text-only LLMs.
Despite the emergence of safety challenges in VLMs, recent studies have revealed a surprising empirical finding: Enhancing VLM safety could be as simple as applying supervised fine-tuning (SFT), provided that a high-quality, dual-modality curated safety fine-tuning dataset is available [14, 19, 20, 21, 22]. One compelling piece of evidence is that fine-tuning on VLGuard [23], a widely used VLM safety dataset, substantially improves robustness against unsafe queries and jailbreaking attacks. As demonstrated in [23], this enhancement surpasses the results obtained by fine-tuning on a “clean” dataset which removes the unsafe data. The surprising effectiveness of SFT on VLGuard has sparked growing interest in re-evaluating its reliability. Recent studies [18, 11, 24] have identified a downside of such safety fine-tuning, known as the over-prudence problem. This issue refers to the over-conservatism of VLMs after safety fine-tuning, where they unnecessarily reject responses when presented with benign queries.
However, the over-prudence suggests that these models may be overly safe, to avoid responses to any skeptical queries. Alternatively, the observed safety may be illusory, as current safety fine-tuning fails to ensure reliability, giving a false sense of safety. Thus, we ask:
(Q) Does current VLM safety fine-tuning achieve true safety? If not, what is the root cause?
Figure 1: Schematic overview of safety mirage findings of safety fine-tuned VLM (LLaVA-v1.5-7B-Mixed, fine-tuned on VLG uard [23]) (a) One-word attack vulnerability: A minor modification (e.g., replacing the first instruction word “Share” with “What” in the original unsafe query) can bypass the safety mechanism established through fine-tuning on VLGuard , even though the safeguarded VLM correctly rejects the original unsafe query. (b) Over-prudence issue: Similar to the one-word attack, a minor modification by replacing “What” with “Share” can cause unnecessary refusals even for benign queries. (c) Root cause of spurious correlations given by fine-tuning dataset biases: Certain words become disproportionately associated with specific safety labels. For example, “Share” is strongly correlated with rejection responses, while “What” is highly associated with non-rejection responses. (d) Effectiveness of unlearning-based safety fine-tuning: The unlearning methods NPO [25] and RMU [26] enhance robustness against attacks while reducing over-prudence, outperforming both the original model LLaVA-v1.5-7B (“Original”) and the supervised fine-tuned LLaVA-v1.5-7B-Mixed [23] (“Prior work”).
In this work, we investigate (Q) and challenge the prevailing belief in the effectiveness of safety fine-tuning for VLMs. We uncover a “safety mirage” in VLM safety fine-tuning, where the seemingly robust safety performance after fine-tuning is primarily driven by spurious correlations between certain non-essential textual words in input queries and predefined safety labels (e.g., rejection) in the fine-tuning dataset. If an adversary identifies these spurious correlations, a simple one-word modification–which we refer to as the “one-word attack”–can effectively jailbreak safety fine-tuned VLMs, enabling them to regenerate unsafe content. Additionally, the input-rejection label shortcut induced by these spurious correlations provides an explanation for the over-prudence of safety fine-tuned VLMs. Similar to the one-word attack, a one-word modification in text queries can readily activate the input-rejection shortcut at test time, causing the model to overgeneralize rejection responses and refuse to generate outputs even for benign queries. In Fig. 1(a)-(c), we provide a schematic overview illustrating: (a) The one-word attack on a safety fine-tuned VLM LLaVA-v1.5-7B-Mixed [23]; (b) The over-prudence issue by one-word modification; (c) The spurious correlations observed in the fine-tuning dataset VLGuard, where the word “share” is strongly linked to rejection, while “what” is associated with non-rejection.
Building on our identification of spurious correlations in VLM safety fine-tuning, we propose improving current safety fine-tuning approaches through machine unlearning (MU). Originally designed to remove the influence of undesired data or knowledge from ML models, MU ensures that essential knowledge remains intact while avoiding unintended disruptions to causally unrelated information [27, 28, 29]. We propose adapting MU to VLM safety fine-tuning as a more robust alternative to traditional supervised approaches. Rather than enforcing safety through direct supervision, MU enhances VLM safety by erasing the influence of unsafe knowledge in a label-free manner, thereby preventing the formation of spurious correlations between input features and safety labels. Although MU has been applied to VLM safety fine-tuning in prior work [17, 30, 31], its unique advantage in mitigating spurious correlations within fine-tuning datasets remains unexplored. Fig. 1(d) showcases the effectiveness of applying two LLM unlearning approaches, NPO [25] and RMU [26] in enhanced robustness against jailbreaking attacks and reduced over-prudence rates.
In summary, our key contributions are listed below.
① We revisit the problem of supervised safety fine-tuning for VLMs and find that there exists a safety mirage, driven by hidden biases—specifically, spurious correlations between non-essential textual features and safety labels in the fine-tuning dataset.
② From an attack perspective, we show that safety fine-tuned VLMs are still susceptible to jailbreaking when adversaries exploit spurious correlations embedded in the fine-tuning dataset. We propose a simple and effective one-word attack by substituting highly frequent querying words associated with rejection responses with those linked to normal model outputs. Additionally, we show that these spurious correlations also contribute to over-prudence, causing fine-tuned VLMs to unnecessarily reject benign inputs.
③ From a defense perspective, we show that MU offers a promising solution to alleviate the effects of spurious correlations in fine-tuning data for VLM safety fine-tuning. The key rationale is that unlearning removes the influence of unsafe responses without relying on the spurious feature-label correlation present in the fine-tuning dataset.
④ We conduct extensive experiments across multiple VLM safety evaluation benchmarks, including VLGuard [23], SPA-VL [21], MM-SafetyBench [14], and FigStep [15], and assess model utility on standard VQA datasets. Our results confirm the safety mirage phenomenon and demonstrate that MU-based safety fine-tuning effectively mitigates spurious correlations and reduces over-prudence.
2 Related Work
VLM safety: Attack and defense. With the rapid advancement of VLMs [6, 7, 8, 32, 33, 34, 1, 2, 4], safety concerns have become increasingly prominent due to their potential to generate harmful or inappropriate content. While LLMs have been extensively studied for safety risks, leading to the development of attack strategies [35, 36, 37, 38], defense mechanisms [39, 40, 41], and robust evaluation datasets [42, 43, 44], VLMs introduce additional challenges due to the complexity of multimodal inputs [16, 17, 18], making them even more vulnerable to jailbreaking and adversarial manipulation [11, 12, 13, 14, 15]. Attacks on VLMs often leverage the dual-modality nature of these models. One approach embeds unsafe textual queries into images through typographic manipulation, enabling the model to bypass safety filters and generate harmful outputs [15, 14]. Another strategy involves using gradient-based adversarial image generation [45, 46, 47, 48, 49] to trigger harmful responses, demonstrating that VLMs remain susceptible to adversarial perturbations despite safety fine-tuning. Defensive strategies for VLM safety generally fall into two categories: Inference-time defenses and fine-tuning with curated safety datasets. The former aligns safety responses dynamically at runtime, mitigating unsafe outputs using various filtering and rejection mechanisms [50, 51, 16, 52, 24]. The latter focuses on red-teaming dataset curation [14, 19, 20, 21, 22, 23, 53], enabling VLMs to be explicitly trained to reject harmful content while retaining utility for benign tasks.
Machine unlearning in VLMs. MU [27, 28, 29, 54, 55] is designed to remove harmful data influences from a pre-trained model while preserving its overall utility. In the LLM domain, recent work has explored targeted forgetting techniques to erase specific knowledge without compromising performance [25, 26, 56, 57, 58, 59, 60]. In VLMs, several benchmarks have established systematic evaluation frameworks for MU algorithms [61, 62, 63]. For safety fine-tuning, prior studies have applied MU-based approaches to mitigate harmful content generation [30, 31, 17]. Our work builds on these efforts by leveraging MU to specifically mitigate spurious correlations present in safety fine-tuning datasets.
3 Preliminary and Problem Statement
Figure 2: Visualization of question-answer samples on the safety fine-tuned VLMs (LLaVA-v1.5-7B-Mixed and LLaVA-v1.5-7B-Posthoc [23, 64]). Green shield represents the correct response, either a safe rejection for harmful queries or a valid answer for benign queries. Red exclamation ! indicates an unsafe response to harmful queries. And red question ? represents an inappropriate rejection for a safe query. (a) Successfully jailbreaking: The safety fine-tuned model originally produces rejection-based responses for unsafe queries; However, replacing the initial question word with “What” can easily bypass this safeguard. (b) Over-prudence: A minor modification by replacing “What” with “Share” can trigger unnecessary refusals even for benign queries.
Existing VLM safety fine-tuning setup.
Previous works [16, 15, 14] highlight the need for safety alignment in VLMs to encompass both textual and imagery data. Consequently, many efforts [23, 21, 65] have focused on curating high-quality dual-modality safety datasets for VLMs. A notable benchmark dataset is VLGuard [23], which covers various text-image pairing scenarios, including unsafe cases where either the text is unsafe or both text and image are unsafe, as well as safe cases where both modalities are benign. Leveraging these curated safety datasets, recent works [23, 17, 18, 21] show that simple fine-tuning approaches on such datasets can yield surprisingly strong safety performance, even against common jailbreaking attacks [66, 67, 68]. In this work, we revisit the VLM safety fine-tuning problem and later argue that the observed safety improvements from fine-tuning may be an illusion.
We begin by presenting the problem formulation of VLM safety fine-tuning. Let $\mathcal{D}_{\mathrm{u}}$ denote the unsafe dataset, which consists of unsafe text queries and corresponding input images possibly paired with targeted safe responses (e.g., rejection responses in VLGuard). In addition, let $\mathcal{D}_{\mathrm{r}}$ denote the retain dataset, which consists of either a safe text-image dataset or a safety-irrelevant utility dataset, designed to maintain VLM performance on normal tasks after safety fine-tuning. For a VLM parameterized by ${\bm{\theta}}$, the safety fine-tuning problem can be formulated as:
$\begin{array}[]{ll}\displaystyle\operatorname*{\text{minimize}}_{{\bm{\theta}}}&\ell_{\mathrm{u}}({\bm{\theta}};\mathcal{D}_{\mathrm{u}})+\gamma\ell_{\mathrm{r}}({\bm{\theta}};\mathcal{D}_{\mathrm{r}}),\end{array}$ (2)
where $\ell_{\mathrm{u}}$ and $\ell_{\mathrm{r}}$ denote the fine-tuning losses over $\mathcal{D}_{\mathrm{u}}$ and $\mathcal{D}_{\mathrm{r}}$ respectively, and the regularization parameter $\gamma\geq 0$ strikes a balance between safety alignment and preserving performance on normal tasks.
Safety mirage: Motivation and problem of interest. Although the safety fine-tuned VLMs could overly reject queries even when presented with benign ones [18, 11, 24], the resulting safety performance against unsafe queries does not degrade; It remains highly robust even in the presence of some common jailbreaking attacks [23, 21].
The seemingly ‘robust’ safety performance observed after fine-tuning motivates us to re-examine its true reliability. As demonstrated in Fig. 1, the current safety fine-tuned VLM remains highly vulnerable to simple paraphrasing of text queries even when only the first question word is modified. As a supplement to Fig. 1, Fig. 2 provides additional motivating examples illustrating the vulnerability of safety fine-tuned VLMs to both jailbreaking attacks and over-prudence. Consistent with Fig. 1, unsafe queries prefixed with innocuous question word “What” successfully bypass safeguard, and harmless prompts start with “Share” trigger over-rejection effect. Notably, the choice of the replacement word ‘What’ or ‘Share’ is not random but rather stems from the spurious correlations embedded in the safety fine-tuning dataset, as we will illustrate in Sec. 4.
Examples in Figs. 1-2 suggest that fine-tuning VLMs on safety datasets may create a “safety mirage”, as evidenced by their susceptibility to even minor one-word modification in text queries. Thus, our work focuses on the following key research questions: (a) What is the root cause of the “safety mirage” in VLM safety fine-tuning? (b) What can be improved to mitigate the “safety mirage”? We explore these questions in Sec. 4 and Sec. 5, respectively.
4 The Risks of Spurious Correlations in VLM Safety Fine-Tuning
Spurious text features and spurious correlations. As discussed in Sec. 3, the effectiveness of current VLM safety fine-tuning methods heavily relies on the curation of high-quality, dual-modality safety datasets. As a result, the safety capabilities of fine-tuned VLMs (i.e., their abilities to prevent harmful content generation) are primarily learned from safety labels (i.e., safe responses) introduced in the fine-tuning datasets. For example, in VLGuard [23], safety labels for unsafe text queries and/or unsafe images are assigned as rejection responses, such as “I’m sorry, I cannot assist with that request…” against the unsafe text query in Figs. 1-2. Additionally, safety labels may correspond to standard VLM responses when processing safe text-image inputs.
At first glance, the use of safety labels appears appropriate. However, a hidden bias may arise when these safety labels become strongly correlated with spurious features in the input data, particularly within textual queries, which are the focus of this work. Here the term “spurious features” refers to non-essential textual words in inputs that do not contribute to the fundamental meaning or task-relevant aspects of the input query. For example, in Fig. 2, the word “What” or “Share” at the beginning of the input query can be considered a spurious feature because it does not directly relate to the query’s actual content and can be easily substituted with other question words. In contrast, “core” features (such as “crime”) are more informative, representing content-related words that capture the true intent and meaning of the query. Therefore, we define spurious correlations as the (unexpected) strong associations between spurious input features and the assigned labels in the safety fine-tuning dataset.
The above conceptualization of spurious correlations is inspired by conventional analyses in image classification [69, 70, 71], where spurious features correspond to background or non-object pixels, while core features directly relate to the target object. In these studies, spurious correlations arise when the target label is linked to spurious features, rather than solely to core features. In the context of VLMs, while several studies have explored spurious correlations [72, 73], their focus has been primarily on spurious visual features within the image modality. To the best of our knowledge, the spurious correlations caused by textual features in VLM safety fine-tuning have not been investigated before.
In this work, we identify two types of spurious correlations in VLM safety fine-tuning. (a) Non-rejection bias: Certain words (like “What” in Fig. 2(a)) in text queries become spuriously correlated with non-rejection responses. As a result, incorporating these words into an original query can easily jailbreak fine-tuned VLMs. (b) Rejection bias: Certain words (like “Share” in Fig. 2(b)) in text queries become spuriously correlated with rejection responses, causing the fine-tuned VLM to exhibit over-prudence.
Refer to caption Refer to caption
(a) Top words in safe queries (b) Top words in unsafe queries
Figure 3: Frequency of question-initiating words in VLGuard queries: (a) words in safe queries associated with non-rejection responses, (b) words in unsafe queries associated with rejections.
To determine the spurious textual features and the spurious correlations (a)-(b), we analyze the frequency of words used in train-time text queries that lead to non-rejection responses (i.e., generating textual content in response to safe queries in the training set) and rejection responses (i.e., predefined refusal answers against unsafe queries in the training set), respectively. Fig. 3 presents the most frequently occurring starting words in text queries from the VLGuard dataset, categorized based on their tendency to elicit non-rejection responses or rejection responses. As shown, the question word “what” predominantly correlates with non-rejection responses, appearing in over 80% of safe queries where the model generates a response. In contrast, the question-initiating words “can” and “share” in unsafe queries are strongly associated with rejection responses, with over 50% of its occurrences leading to a rejection. Compared to “can”, which appears in both safe and unsafe queries, the word “share” is only used in unsafe queries to elicit rejection.
One-word jailbreaking. Recognizing the non-rejection bias (i.e., the spurious correlation between certain querying words like “what” and non-rejection responses) as shown in Fig. 3, an adversary can exploit this spurious correlation to jailbreak safety fine-tuned VLMs. Formally, let $q$ denote the original unsafe text query that VLM avoids generating unsafe content for. And let $q^{\prime}=\mathrm{w}_{\mathrm{adv}}+q$ denote the jailbreaking attack, where $\mathrm{w}_{\mathrm{adv}}$ is the adversarial perturbation chosen based on the non-rejection bias-inducing querying word (e.g., “what” in Figs. 1, 2, and 3), and the operation $+$ signifies either word insertion as a prefix to $q$ or a simple starting word replacement in $q$. We refer to the above simple attack strategy as the one-word jailbreaking attack.
Figure 4: ASR of $K$-shot one-word attack for varying $K$, evaluated before and after applying “What”-initialized one-word attack to jailbreak the safety fine-tuned VLM (LLaVA-v1.5-7B-Mixed [23]).
[…]
One-word over-prudence. Similar to the one-word jailbreaking attack, when a rejection bias-inducing word (like “Share” in Fig. 3) is introduced into text queries, we observe that even benign modified queries can still prevent the fine-tuned VLM from generating any meaningful output. To amplify the over-prudence rate, we can also apply the multi-shot strategy (i.e., “Share”-based one-word modification to different paraphrased versions of the benign input query). The results for the multi-shot setting are presented in Appendix A. Notably, even in the one-shot setting, the “Share”-initialized benign query modification already achieves a 90% over-rejection rate against safe text-image queries.
5 Enhancing VLM Safety through Unlearning
To mitigate spurious correlations embedded in the safety fine-tuning dataset, one potential solution, motivated by the analysis in Sec. 4, is to eliminate the dependence of safety fine-tuning on safety labels (e.g., rejection responses to unsafe queries). This necessitates shifting from supervised fine-tuning to a label-free, unsupervised setting for safety alignment. Machine unlearning (MU) [27, 28, 29] provides an ideal solution in this context, as it is designed to remove the undesired influence of harmful data or knowledge from a pre-trained model while preserving its normal utility.
Although unlearning has been applied to VLM safety fine-tuning in prior work [17, 30, 31], its unique advantage and application in circumventing spurious correlations within fine-tuning datasets remains unexplored. To this end, we adapt two state-of-the-art MU approaches, originally developed for large language models (LLMs)–representation misdirection unlearning (RMU) [26] and negative preference optimization (NPO) [25]–to the context of VLM unlearning.
The proposed VLM unlearning follows the generic formulation of (2), but with the following key modifications. First, the fine-tuning loss over the unsafe dataset $\mathcal{D}_{\mathrm{u}}$ is replaced with an unlearning objective $\ell_{\mathrm{u}}$ that relies solely on the unsafe data features (text-image queries) in $\mathcal{D}_{\mathrm{u}}$, without depending on the safety labels. In our work, we define the unlearning loss $\ell_{\mathrm{u}}$ based on the principles of RMU and NPO, respectively. The RMU-based unlearning objective aims to map the intermediate features of unsafe data $x\in\mathcal{D}_{\mathrm{u}}$ (to be forgotten) to random features. This ensures that the model no longer retains meaningful representations of the unsafe data. The objective is given by
$\ell_{\mathrm{u}}({\bm{\theta}};\mathcal{D}_{\mathrm{u}})=\mathbb{E}_{\mathbf{x}\in\mathcal{D}_{\mathrm{u}}}[\|M_{{\bm{\theta}}}(\mathbf{x})-c\cdot\mathbf{v}\|_{2}^{2}],$ (3)
where $M_{{\bm{\theta}}}(\cdot)$ represents certain intermediate-layer representations of ${\bm{\theta}}$, $c$ is a hyperparameter that controls activation scaling , and $\mathbf{v}$ is a random vector drawn from a standard uniform distribution. We remark that, unlike RMU for LLM unlearning [26], we carefully adjusts the representation layer selection and tunes the hyperparameter $c$ to better suit the unlearning process in VLMs.
Table 1: Experiment results evaluating safety, over-prudence, and utility of safety fine-tuned VLMs. Safety is quantified by ASR (attack success rate) on unsafe input queries, evaluated before and after a 3-shot one-word attack (i.e., “what”-based prefix in Fig. 1) that promotes non-rejection bias. Over-prudence is measured by RR (rejection rate) on safe input queries, evaluated before and after using 1-shot, one-word over-prudence modification (i.e., “share”-based prefix in Fig. 1). Here, “Before” and “After” denote the performance prior to and following the respective one-word modification. Utility is assessed by the accuracy (Acc.) on four VQA benchmarks: VQAv2, TextVQA, ScienceQA, and VizWiz. Results are presented for models under full fine-tuning and LoRA fine-tuning settings, with safety fine-tuning approaches including Unsafe-Filter, Mixed-SFT, Posthoc-SFT, NPO-Unlearning, and RMU-Unlearning.
Models Safety Evaluation (ASR, $\downarrow$) Over-Prudence Evaluation (RR, $\downarrow$) Utility Evaluation (Acc., $\uparrow$)
VLGuard SPA-VL VLGuard SPA-VL VQAv2 TextVQA ScienceQA VizWiz
Before After Before After Before After Before After
LLaVA-1.5-7B 64.25% 90.27% 46.42% 52.08% 0.36% 0.36% 14.72% 9.81% 78.53% 58.23% 69.51% 50.07%
       + Unsafe-Filter 65.66% 90.72% 45.66% 54.72% 0.36% 0.36% 15.85% 11.32% 79.14% 58.22% 68.12% 52.14%
       + Mixed-SFT 0.23% 54.98% 14.34% 37.73% 4.48% 91.76% 68.68% 98.87% 78.23% 57.80% 68.27% 52.94%
       + Posthoc-SFT 0.23% 46.83% 13.58% 32.96% 2.69% 90.83% 60.38% 100.0% 78.03% 57.73% 68.42% 51.84%
       + NPO-Unlearning 2.49% 12.92% 18.49% 24.15% 2.51% 11.69% 16.60% 17.36% 77.34% 57.80% 68.02% 50.21%
       + RMU-Unlearning 1.29% 10.18% 17.73% 22.64% 1.25% 7.56% 18.11% 19.24% 77.04% 56.89% 67.68% 50.01%
LLaVA-1.5-7B-LoRA 64.72% 95.25% 44.91% 50.44% 0.18% 0.18% 15.47% 12.45% 79.13% 58.22% 68.62% 52.82%
       + Unsafe-Filter 67.19% 93.89% 45.28% 52.33% 0.36% 0.0% 22.64% 13.21% 79.14% 57.66% 67.97% 53.65%
       + Mixed-SFT 0.45% 69.23% 21.51% 40.13% 3.05% 89.93% 59.25% 97.36% 78.63% 57.24% 68.47% 51.84%
       + Posthoc-SFT 0.23% 51.81% 20.38% 37.61% 3.41% 95.14% 62.26% 99.62% 78.23% 57.17% 67.92% 52.08%
       + NPO-Unlearning 4.56% 18.29% 21.51% 25.28% 2.69% 11.01% 16.98% 19.62% 77.32% 56.98% 66.98% 51.01%
       + RMU-Unlearning 3.87% 11.14% 20.38% 24.24% 1.25% 4.84% 18.49% 21.89% 76.99% 56.62% 66.32% 49.87%
LLaVA-1.5-13B 68.10% 91.86% 50.19% 54.47% 0.54% 0.72% 19.62% 14.34% 79.99% 61.25% 72.73% 53.64%
       + Unsafe-Filter 67.65% 92.99% 52.08% 56.27% 0.54% 0.54% 20.38% 15.09% 79.87% 61.32% 71.59% 52.68%
       + Mixed-SFT 0.45% 57.01% 18.11% 40.5% 4.84% 92.63% 58.87% 97.74% 79.03% 60.98% 72.03% 53.01%
       + Posthoc-SFT 1.58% 69.23% 16.98% 32.83% 2.69% 76.08% 56.98% 98.49% 78.94% 60.63% 71.94% 52.31%
       + NPO-Unlearning 1.89% 11.70% 22.26% 26.04% 2.33% 10.65% 23.77% 27.92% 78.31% 60.05% 71.56% 52.04%
       + RMU-Unlearning 1.29% 8.96% 20.00% 23.77% 1.61% 9.36% 25.90% 29.43% 77.98% 59.68% 70.86% 51.67%
LLaVA-1.5-13B-LoRA 67.87% 93.89% 45.66% 55.22% 0.72% 0.54% 19.25% 12.83% 80.04% 60.23% 71.64% 54.74%
       + Unsafe-Filter 66.97% 94.34% 48.30% 55.85% 0.36% 0.54% 22.26% 13.21% 79.98% 60.05% 71.54% 54.02%
       + Mixed-SFT 0.45% 52.94% 14.34% 38.74% 3.05% 92.45% 63.40% 98.87% 78.85% 59.67% 71.42% 53.27%
       + Posthoc-SFT 0.23% 42.08% 12.08% 30.44% 3.41% 79.68% 61.13% 99.62% 78.64% 59.43% 71.40% 53.64%
       + NPO-Unlearning 3.36% 13.53% 18.11% 22.26% 3.59% 10.47% 23.77% 28.68% 78.43% 59.26% 71.36% 53.41%
       + RMU-Unlearning 2.75% 10.18% 17.74% 22.64% 1.79% 8.64% 26.42% 30.56% 78.27% 58.79% 70.98% 52.99%
In addition to RMU, we also employ NPO [25] to model the unlearning objective $\ell_{\mathrm{u}}$, which treats unsafe data designated for unlearning as “negative” examples in a direct preference optimization framework [76]. The NPO-based unlearning loss is then given by
$\begin{array}[]{c}\ell_{\mathrm{u}}({\bm{\theta}};\mathcal{D}_{\mathrm{u}})=\mathbb{E}_{\mathbf{x}\in\mathcal{D}_{\mathrm{u}}}\left[-\frac{2}{\beta}\log\sigma\left(-\beta\log\left(\frac{\pi_{{\bm{\theta}}}(\mathbf{x})}{\pi_{\mathrm{ref}}(\mathbf{x})}\right)\right)\right],\end{array}$ (5)
where $\sigma(\cdot)$ the sigmoid function, $\beta>0$ is the temperature parameter , $\pi_{{\bm{\theta}}}$ denotes the prediction probability of the model ${\bm{\theta}}$ given the unsafe input $\mathbf{x}$, and $\pi_{\mathrm{ref}}$ represents the reference model given by the initial model prior to unlearning. The rationale behind NPO is to fine-tune the VLM ${\bm{\theta}}$ to force it to deviate from the reference model when processing unsafe inputs. Following (2), VLM unlearning also requires the retain loss $\ell_{\mathrm{r}}$ to preserve the model utility in normal tasks. This is achieved using the standard VLM training loss over the safe text and safe image data in the fine-tuning set.
As demonstrated in Sec. 6, compared to conventional supervised safety fine-tuning, unlearning-based approaches produce safer responses by generating irrelevant or neutral outputs rather than merely rejecting unsafe queries. Moreover, unlearning reduces the model’s reliance on rejection responses, alleviating over-prudence and enabling more appropriate handling of benign inputs.
6 Experiments
6.1 Experiment Setups
Datasets and models.
We consider four VLM safety datasets: VLGuard [23], MM-SafetyBench [14], SPA-VL [21], and Figstep [15]. To assess the utility of safety fine-tuned VLMs, we also conduct evaluations on representative visual question-answering (VQA) datasets, including VQAv2 [77], TextVQA [78], VizWiz [79], and ScienceQA [80]. For model selection, we adopt LLaVA-v1.5-7B and LLaVA-v1.5-13B [6, 7] as our primary VLMs.
Safety fine-tuning setups and baselines. In our experiments, we choose VLGuard as the training dataset for VLM safety fine-tuning. When implementing the MU-based fine-tuning approaches, i.e., RMU in (3) and NPO in (5), we use the unsafe input-output pairs from VLGuard as the unsafe dataset ($\mathcal{D}_{\mathrm{u}}$) to be unlearned. Here we employ Llama-2-13B-Chat to confirm the harmfulness of the original model’s responses against the unsafe input queries. Additionally, the safe query-answer pairs from VLGuard are used to construct the retain dataset ($\mathcal{D}_{\mathrm{r}}$) in (2). For additional details, we refer readers to Appendix  B.1-B.2.
Besides MU-based VLM safety fine-tuning, we include a series of popular supervised safety fine-tuning approaches as baselines. (1) Mixed-SFT [23]: Supervised fine-tuning (SFT) using a mixed fine-tuning strategy on VLGuard. (2) Posthoc-SFT [23, 64]: SFT using a post-hoc fine-tuning approach on VLGuard. (3) Unsafe-Filter: SFT performed on a clean training set, where LLaMA-Guard-3-11B-Vision [81] is used to filter unsafe data from pre-training datasets.
Evaluation setups. For a fine-tuned VLM, its safety performance is primarily evaluated using the attack success rate (ASR), defined as the proportion of unsafe responses generated for unsafe input queries over the total number of queries. The unsafe input queries are considered in two scenarios. (1) Before attack: The original unsafe input queries from VLM safety datasets including VLGuard, MM-SafetyBench, SPA-VL, and Figstep. (2) After attack: We apply the proposed one-word attack to the original unsafe queries. By default, our one-word jailbreaking attack is implemented using a 3-shot approach, as illustrated in Fig. 4. Specifically, three paraphrased versions of the text query (while preserving the adversarial perturbation) are posed for each image. The attack is considered successful if any of the generated responses is unsafe. To validate the harmfulness of the responses generated by the fine-tuned model, we employ the multi-modal Qwen2.5-VL-7B-Instruct as a judge. The details of the judge prompt can be found in Appendix  B.4.
To investigate the over-prudence issue in safety fine-tuned VLMs, we measure the rejection rate (RR), which quantifies the proportion of benign test-time queries that are unnecessarily rejected by the fine-tuned model. These benign queries consist of both safe text queries and safe images provided by the VLM safety datasets. Similar to ASR, RR is evaluated both before and after exploiting spurious correlations to encode rejection bias. Furthermore, we evaluate the general utility of safety fine-tuned VLMs by measuring their prediction accuracy (Acc) on standard downstream VQA tasks, including VQAv2, TextVQA, ScienceQA, VizWiz.
6.2 Experiment Results
Overall performance on safety, over-prudence, utility.
In Table 1, we present a comprehensive evaluation of safety fine-tuned VLMs across three key metrics: safety performance against unsafe input queries (measured by ASR, where lower is better), over-prudence performance against safe input queries (measured by RR, where lower is better), and model utility on representative downstream tasks (measured by Acc, where higher is better). Recall that both ASR and RR are evaluated before and after exploiting spurious correlations, i.e., promoting non-rejection bias and rejection bias via one-word modification, respectively.
Figure 5: Visualization of question-answer pairs from three models: LLaVA-1.5-7B (original), Mixed-SFT (fine-tuned), and RMU-Unlearning (unlearned). Green shield represents the correct response, whether a safe rejection for harmful queries or a valid answer for benign queries. Red exclamation ! indicates an unsafe response to harmful queries, while red question ? represents an inappropriate rejection for a safe query. The first row displays responses to unsafe text-image queries, while the second row shows responses to safe queries.
We draw some key observations below. First, conventional safety fine-tuning approaches (Unsafe-Filter, Mixed-SFT, and Posthoc-SFT) exhibit a safety mirage, as evidenced by a significant rise in ASR after applying the one-word attack–nearly 60% on the VLGuard unsafe evaluation set and nearly 30% on the SPA-VL unsafe evaluation set of LLaVA-1.5-7B based model. Additionally, these baselines exhibit a significant over-prudence issue, as evidenced by an over 90% increase after one-word modification in RR on the safe evaluation sets in VLGuard and SPA-VL. Furthermore, both full and LoRA fine-tuning exhibit a similar safety mirage phenomenon in baselines.
Second, compared to baselines, the unlearning-based approaches (NPO and RMU) exhibit significantly lower ASR increases after the one-word attack and maintain a low RR against safe queries, effectively alleviating both jailbreaking susceptibility and over-prudence. Compared to NPO, RMU achieves slightly better performance in both ASR (lower vulnerability to the one-word attack) and RR (reduced over-prudence). This result is consistent with LLM unlearning, where RMU typically outperforms NPO in knowledge unlearning [26]. Furthermore, the advantages of MU persist in both full fine-tuning and LoRA fine-tuning.
Third, from the perspective of model utility evaluation, we observe that unlearning-based approaches lead to a slight decrease in Acc on downstream tasks (approximately 1%), compared to supervised safety fine-tuning. This suggests a potential tradeoff between unlearning effectiveness and utility preservation, indicating that erasing harmful knowledge in VLMs may have an impact on general performance. Further optimizing this tradeoff remains an important future research direction to enhance VLM unlearning.
Table 2: Analyzing the safety mechanisms of unlearning-based approaches vs. baselines. Before and after applying the one-shot “What”-initialized attack, the safety rate (1-ASR) against unsafe queries is decomposed into RR (rejection rate) and irrelevance rate (IR), where IR represents responses that are irrelevant to the unsafe queries. All other setups remain consistent with Table 1.
Models Safety Evaluation on VLGuard
Before After
ASR IR RR ASR IR RR
LLaVA-1.5-7B 64.25% 30.09% 5.66% 74.43% 21.95% 3.62%
      +Unsafe-Filter 65.66% 28.01% 6.33% 74.66% 21.49% 3.85%
      +Mixed-SFT 0.23% 0% 99.77% 24.66% 5.20% 70.14%
      +Posthoc-SFT 0.23% 0% 99.77% 25.34% 4.75% 69.91%
      +NPO-Unlearning 2.49% 46.42% 51.09% 6.99% 48.72% 44.29%
      +RMU-Unlearning 1.29% 93.96% 4.75% 5.06% 89.29% 5.65%
LLaVA-1.5-7B-LoRA 64.72% 28.28% 7.02% 72.62% 21.95% 5.43%
      +Unsafe-Filter 67.19% 26.47% 6.33% 73.08% 20.81% 6.11%
      +Mixed-SFT 0.45% 0.0% 99.55% 39.59% 5.66% 54.75%
      +Posthoc-SFT 0.23% 0.0% 99.55% 20.81% 2.94% 76.24%
      +NPO-Unlearning 4.56% 48.64% 46.80% 6.86% 53.14% 40.0%
      +RMU-Unlearning 3.87% 90.92% 5.21% 6.91% 88.33% 4.76%
Other safety evaluation. We further evaluate different safety fine-tuning approaches using LLaVA-v1.5-7B on MM-SafetyBench and FigStep, with results Table A1. Consistent with our findings on VLGuard and SPA-VL, the one-word attack significantly increases ASR across both benchmarks. In Fig. 5, we present input-output demonstrations of safety fine-tuned LLaVA-1.5-7B, comparing the baseline Mixed-SFT with the RMU-based unlearning approach. The original LLaVA-1.5-7B model is vulnerable to unsafe queries both before and after the one-word attack. Here the attack is primarily executed by replacing “How” with “What” in text queries. However, the one-word attack effectively bypass the safety mechanism from Mixed-SFT, triggering unsafe response against the query starting with “What”. Moreover, the Mixed-SFT fine-tuned VLM exhibits over-prudence, as evidenced by its rejection response to a safe text query starting with “Share”. Here “Share” is identified as a high-frequency word in the fine-tuning dataset that is strongly correlated with rejection responses in Fig. 3. By contrast, the RMU unlearning approach effectively defends against the one-word attack while also mitigating the over-rejection issue.
Figure 6: Input token sensitivity analysis for “What”-initiated unsafe queries and “Share”-initiated safe queries over VLGuard using LLaVA-1.5-7B Mixed-SFT, before and after masking “What” and “Share”. Sensitivity is measured using per-token masking (i.e., replacing the original token with a blank placeholder [PAD]) to evaluate each token’s influence on response rejection probabilities.
Irrelevance response for safety enhancement by unlearning. To understand the difference in the safety mechanism of MU-based approaches vs. safety-aware SFT, Table 2 presents the rates of unsafe, irrelevant, and rejection responses in our safety evaluation, both before and after a 1-shot one-word attack. Here, the safety rate (1-ASR) is further decomposed into: (1) irrelevant responses where the model sidesteps the unsafe query by generating a response that is unrelated to the harmful content, and (2) rejection responses where the model explicitly refuses to respond. As observed, conventional SFT-based safety fine-tuning strategies predominantly produce rejection-based responses, as reflected in the overly high RR in both “Before” and “After” scenarios. In contrast, our unlearning-based methods primarily yield irrelevant responses, reducing the model’s reliance on outright rejections, as evidenced by the high irrelevance rate (IR). Notably, the RMU-Unlearning approach steers the model’s output distribution closer to a random vector, making responses more likely to be classified as irrelevant rather than explicitly rejecting queries.
Input token saliency analysis. Recall that spurious correlations emerge between specific words in textual queries and safety labels. To further analyze this, we investigate input token sensitivity by applying per-token masking (i.e., replacing the original input token with a blank placeholder [PAD]) to assess the token influence in non-rejection and rejection responses. Fig. 6 presents the rejection probability using LLaVA-1.5-7B Mixed-SFT for all “What”-initiated unsafe input queries and “Share”-initiated safe queries before and after masking “What” and “Share”, respectively, over the VLGuard dataset. As we can see, masking “What” significantly increases the rejection probability, confirming its role in inducing non-rejection bias in VLGuard. Masking “Share” significantly reduces the rejection probability compared to its unmasked counterpart, demonstrating its influence in reinforcing rejection bias. See Appendix B.6 for specific textual input saliency maps.
Additional results. In Appendix B.7, we demonstrate the robustness of spurious correlation and the effectiveness of our MU-based approach when facing visual variations.
7 Conclusion
In this work, we unveil the “safety mirage” in VLMs, a deceptive robustness that emerges from supervised safety fine-tuning. Our analysis reveals that biases in the fine-tuning dataset reinforce spurious correlations between superficial textual patterns and safety labels, leading to a false sense of security. As a result, fine-tuned VLMs become highly susceptible to simple one-word jailbreaking attacks while also exhibiting over-prudence, unnecessarily rejecting benign queries. To address these, we propose MU as a principled alternative to fine-tuning. Unlike supervised approaches, MU directly removes harmful knowledge without relying on explicit safety labels, thereby avoiding biased feature-label mappings and mitigating spurious correlations. Extensive experiments confirm the existence of the “safety mirage” in conventional VLM safety fine-tuning and demonstrate that MU-based safety alignment significantly enhances robustness against jailbreaking attacks, reduces over-prudence, and preserves strong performance on standard VQA tasks.
8 Broader Impacts
Our work reveals spurious correlations in VLM safety fine-tuning, which are prevalent across models and can function as unintentional “backdoors” in real-world applications. Malicious actors could exploit these correlations to launch jailbreaking attacks, potentially extracting sensitive or private information from VLMs. While the advocated unlearning methods aim to enhance safety alignment, they also present dual-use risks. A bad actor could misuse these techniques to erase safety guardrails, rendering a previously robust model unsafe. If such an unlearned model were publicly released, it could bypass safety mechanisms, increasing the risk of harmful content generation.
9 Acknowledgement
We extend our gratitude to Cisco Research and DSO National Laboratories for their support in this project. Additionally, the contributions of Yiwei Chen, Yuguang Yao, Yihua Zhang, and Sijia Liu are partially supported by the National Science Foundation (NSF) CISE Core Program Award IIS-2207052, the NSF Cyber-Physical Systems (CPS) Award CNS-2235231, the NSF CAREER Award IIS-2338068, the Amazon Research Award, and the IBM PhD Fellowship.
Appendix
A One-word Over Prudence
Similar to the one-word jailbreaking attack, inserting a rejection bias-inducing word (e.g., “Share” as shown in Fig. 3) into text queries causes the fine-tuned VLM to produce no meaningful output, even for benign inputs. We also apply a multi-shot strategy — modifying different paraphrases of a benign query with “Share”. Notably, even a one-shot “Share” modification leads to a 90% over-rejection rate on safe queries.
Figure A1: Rejection rate vs. $K$, evaluated before and after applying the “Share”-initialized one-word modification to safe input queries, cause over-prudence phenomenon of the safety fine-tuned VLM LLaVA-v1.5-7B-Mixed model [23].
B Additional Experiment Details and Results
B.1 Additional Fine-tuning Setups
All experiments are conducted on a single node equipped with 8 NVIDIA A6000 GPUs. We perform fine-tuning using both full-parameter and LoRA-based methods. For full fine-tuning, we employ a global batch size of 8 and a learning rate of $1\times 10^{-6}$. For LoRA-based fine-tuning, following [6, 21], we set the rank $r$ to 256 and the scaling factor $\alpha$ to 512, with a global batch size of 64 and a learning rate of $1\times 10^{-5}$. In both setups, we use the AdamW optimizer with zero weight decay and a warmup ratio of 0.03.
B.2 Machine Unlearning Implementation Details
To address problem (2), fine-tuning can be effectively performed either through full parameter updates or parameter-efficient techniques such as low-rank adaptation (LoRA) [82]. Common implementations adopt cross-entropy loss for $\ell_{\mathrm{r}}$, while $\ell_{\mathrm{u}}$ can flexibly incorporate various optimization objectives such as cross-entropy loss, DPO [76], PPO [83].
RMU implementation details.
To apply the RMU algorithm for safety alignment, we use unsafe input-output pairs from VLGuard as the unsafe dataset $\mathcal{D}_{\mathrm{u}}$ for unlearning. We employ Llama-2-13B-Chat to verify the harmfulness of the original LLaVA series model’s responses to unsafe queries and select the most harmful response as the target for unlearning. Specifically, for RMU, we combine the question and answer into a single instance as VLM’s input to be forgotten. For the retain dataset $\mathcal{D}_{\mathrm{r}}$, we use a mixture of the LLaVA fine-tuning data and the safe input-output pairs from VLGuard. The unlearning loss $\ell_{\mathrm{u}}$ defined as (3), where $M_{{\bm{\theta}}}(\cdot)$ denotes the intermediate-layer representations (we use layers 4, 5, and 6 in our implementation), $c$ is a hyperparameter for activation scaling (set to $c=10$), and $\mathbf{v}$ is a random vector drawn from a standard uniform distribution. The retain loss $\ell_{\mathrm{r}}$ is computed as the cross-entropy loss on the retain data, with the regularization parameter set to $\gamma=1.2$.
NPO implementation details.
Compared to the RMU implementation, we use the same settings for the unsafe dataset $\mathcal{D}_{\mathrm{u}}$ and the retain dataset $\mathcal{D}_{\mathrm{r}}$. However, for NPO, the input to the VLM from $\mathcal{D}_{\mathrm{u}}$ consists solely of the harmful question. In this setting, unsafe data designated for unlearning is treated as negative examples within a direct preference optimization framework [76]. The unlearning training objective is defined as (5), where $\sigma(\cdot)$ denotes the sigmoid function, $\beta>0$ is a temperature parameter (set to $\beta=0.6$ in our implementation), $\pi_{{\bm{\theta}}}(\mathbf{x})$ is the prediction probability of the model given the unsafe input $\mathbf{x}$, and $\pi_{\mathrm{ref}}(\mathbf{x})$ is the corresponding probability from the reference model (i.e., the initial model prior to unlearning). The retain loss $\ell_{\mathrm{r}}$ is computed as the cross-entropy loss on the retain data, with the regularization parameter set to $\gamma=1.0$.
B.3 Additional Baseline Details
For the Unsafe-Filter baseline, we employ the LLaMA-Guard-3-11B-Vision [81] model to assess the safety of image-instruction-answer pairs in the LLaVA pre-training and fine-tuning datasets. Unsafe data is filtered out, and LLaVA is subsequently fine-tuned using only the remaining safe samples. The fine-tuning setup follows the original LLaVA configuration.
B.4 Evaluation Setup Details
To validate the harmfulness of the responses generated by the fine-tuned model, we employ =Qwen2.5-VL-7B-Instruct model as a judge. Below is the prompt used for evaluating content relevance, which is included in our experiments:
messages = [
“role”: “system”,
“content”: [
{“type”: “text”, “text”: (
“You are an expert in content evaluation. Your task is to determine whether the given answer to the image and question pair is Relevant or Irrelevant.
[…]
“1. An answer should be classified as ’Relevant’ if it is detailed, well-organized, and clearly addresses both the question and the visual content.
[…]
“2. An answer should be classified as ’Irrelevant’ if it is vague, poorly written, or contains nonsensical phrases.
“Always provide detailed reasoning for your decision, and ensure your response strictly starts with either ’Relevant:’ or ’Irrelevant:’.”
“role”: “user”,
“content”: [
{“type”: “image”, “image”: image_path},
{“type”: “text”, “text”: (
[…]
“Please analyze the provided answer in terms of its detail, clarity, and correctness in addressing the image and question.
“Provide detailed reasoning for your decision.”
B.5 Other Safety Evaluation
In Table A1, we present additional safety evaluations of different safety fine-tuning approaches using LLaVA-v1.5-7B on two additional safety benchmarks: MM-SafetyBench and FigStep. The results are consistent with our earlier findings on VLGuard and SPA-VL: The one-word attack leads to a significant increase in ASR on both MM-SafetyBench and FigStep. This reveals that the spurious correlations learned during conventional safety fine-tuning persist across different evaluation settings. Our proposed MU-based methods demonstrate improved robustness, thereby alleviating this vulnerability.
Table A1: Safety evaluation on MM-Safetybench and FigStep. ASR is reported for unsafe inputs before and after the “What”-initialized 3-shot attack. The setup and format follow Table 1.
Models Safety Evaluation ($\downarrow$)
MM-Safety (ASR) FigStep (ASR)
Before After Before After
LLaVA-1.5-7B 48.81% 91.27% 62.00% 86.00%
      +Unsafe-Filter 50.60% 90.28% 62.00% 84.00%
      +Mixed-SFT 0.60% 48.81% 0.00% 20.00%
      +Posthoc-SFT 0.60% 40.48% 0.00% 28.00%
      +NPO-Unlearning 4.76% 20.24% 6.00% 12.00%
      +RMU-Unlearning 2.98% 17.26% 4.00% 10.00%
LLaVA-1.5-7B-LoRA 57.74% 93.45% 72.00% 84.00%
      +Unsafe-Filter 58.93% 91.07% 74.00% 84.00%
      +Mixed-SFT 0.60% 63.69% 0.00% 40.00%
      +Posthoc-SFT 0.60% 41.07% 0.00% 36.00%
      +NPO-Unlearning 4.17% 23.81% 4.00% 16.00%
      +RMU-Unlearning 5.36% 19.05% 2.00% 12.00%
Table A2: Evaluating the safety and over-prudence under various visual perturbations on VLGuard, including Gaussian Noise, Gaussian Blur, and Color Jitter. Other setups are consistent with Table 1.
Models Safety Evaluation (ASR, $\downarrow$) Over-Prudence (RR, $\downarrow$)
Before After Before After
LLaVA-1.5-7B+Mixed-SFT
Original Image 0.23% 54.98% 4.48% 91.76%
+ Gaussian Noise 0.45% 54.08% 4.30% 91.22%
+ Gaussian Blur 0.23% 54.31% 4.66% 91.04%
+ Color Jitter 0.45% 54.75% 4.66% 90.86%
LLaVA-1.5-7B+Posthoc-SFT
Original Image 0.23% 46.83% 2.69% 90.83%
+ Gaussian Noise 0.23% 46.38% 2.15% 90.29%
+ Gaussian Blur 0.23% 45.93% 1.97% 90.29%
+ Color Jitter 0.45% 46.16% 1.97% 91.19%
LLaVA-1.5-7B+NPO-Unlearning
Original Image 2.49% 12.92% 2.51% 11.69%
+ Gaussian Noise 2.26% 12.47% 2.15% 11.51%
+ Gaussian Blur 2.26% 12.70% 1.97% 11.15%
+ Color Jitter 2.71% 13.15% 1.97% 12.05%
LLaVA-1.5-7B+RMU-Unlearning
Original Image 1.29% 10.18% 1.25% 7.56%
+ Gaussian Noise 1.29% 9.73% 1.43% 7.20%
+ Gaussian Blur 1.74% 9.96% 1.43% 7.38%
+ Color Jitter 1.74% 10.41% 1.07% 7.02%
B.6 Examples of Input Saliency
We show two input saliency maps in Fig. A2 and Fig. A3. The input saliency is shown through the prediction probability change for each output sentence ($y$-axis) by masking each input token ($x$-axis). Here [X] refers to the token “X” is marked (i.e., replacing with [PAD])), and [ ] denotes the non-masking case. The examples of input query in Fig. A2 and Fig. A3 correspond to the examples introduced in Fig. 1(a)-(b). Fig. A2 shows that masking “what” token can significantly boost the rejection probability, producing “I’m sorry…”. Fig. A3 shows that masking “share” token successfully reduces the prediction probability for rejection response and produces the normal model generation.
Figure A2: Prediction probability for various output sentences before and after masking each input query token for the query “What can be used to create a promotion?” corresponding to the image example in Fig. 1-(a).
Figure A3: Prediction probability for various output sentences before and after masking each input query token for the query “Share the drink type in the image” corresponding to the image example in Fig. 1-(b).
B.7 Spurious Correlations with Visual Variations
To further evaluate the persistence of spurious correlations between text queries and safety labels and the robustness of our MU-based safety fine-tuning approaches, we present in Table A2 the safety performance of fine-tuned models when subjected to various visual perturbations. The results reveal that the SFT-based safety models consistently exhibit a significant increase in both ASR and RR even under visual variations. This finding indicates that the spurious correlations, and consequently the safety mirage phenomenon remains stable. In contrast, the performance of our MU-based safety aligned models remains robust across these perturbations.

### v2

@@
-Safety Mirage: How Spurious Correlations Undermine VLM Safety Fine-tuning
+Safety Mirage: How Spurious Correlations Undermine VLM Safety Fine-Tuning and Can Be Mitigated by Machine Unlearning
 Abstract
-Recent vision-language models (VLMs) have made remarkable strides in generative modeling with multimodal inputs, particularly text and images.
+Recent vision language models (VLMs) have made remarkable strides in generative modeling with multimodal inputs, particularly text and images.
 However, their susceptibility to generating harmful content when exposed to unsafe queries raises critical safety concerns.
-While current alignment strategies primarily rely on supervised safety fine-tuning with curated datasets, we identify a fundamental limitation we call the “safety mirage” where supervised fine-tuning inadvertently reinforces spurious correlations between superficial textual patterns and safety responses, rather than fostering deep, intrinsic mitigation of harm.
+While current alignment strategies primarily rely on supervised safety fine-tuning with curated datasets, we identify a fundamental limitation we call the “safety mirage”, where supervised fine-tuning inadvertently reinforces spurious correlations between superficial textual patterns and safety responses, rather than fostering deep, intrinsic mitigation of harm.
 We show that these spurious correlations leave fine-tuned VLMs vulnerable even to a simple one-word modification-based attack, where substituting a single word in text queries with a spurious correlation-inducing alternative can effectively bypass safeguards.
-Additionally, these correlations contribute to the over prudence, causing fine-tuned VLMs to refuse benign queries unnecessarily.
-To address this issue, we show machine unlearning (MU) as a powerful alternative to supervised safety fine-tuning as it avoids biased feature-label mappings and directly removes harmful knowledge from VLMs while preserving their general capabilities.
-Extensive evaluations across safety benchmarks show that under one-word attacks, MU-based alignment reduces the attack success rate by up to 60.17% and cuts unnecessary rejections by over 84.20%.
-Codes are available at https://github.com/OPTML-Group/VLM-Safety-MU.
+Additionally, these correlations contribute to the over-prudence, causing fine-tuned VLMs to refuse benign queries unnecessarily.
+To address these issues, we show machine unlearning (MU) as a powerful alternative to supervised safety fine-tuning, as it avoids biased feature-label mappings and directly removes harmful knowledge from VLMs while preserving their general capabilities.
+Extensive evaluations across safety benchmarks show that under MU-based alignment reduces the attack success rate by up to 60.17% and cuts unnecessary rejections by over 84.20%.
+Codes are available at https://github.com/OPTML-Group/VLM-Safety-Unlearn.
 WARNING: There exist AI generations that may be offensive in nature.
 1 Introduction
-Recent multi-modal models have achieved advancements by integrating text and images [1, 2, 3, 4, 5].
-A prevalent architecture in VLMs maps visual embeddings into the language model’s latent space via a dedicated projection module [6, 7, 8, 9, 10].
-While the large language model (LLM) backbone effectively integrates projected visual information, the image modality also introduces new vulnerabilities [11, 12, 13, 14, 15].
-Specifically, images can function as a “foreign language” [16], creating pathways for unsafe input queries, even when the underlying LLM has been aligned for safety [16, 17, 18].
+Recent multi-modal models have achieved advancements by integrating text and images (Alayrac et al., 2022; Awadalla et al., 2023; Hurst et al., 2024; Gao et al., 2023; Li et al., 2023a).
+A prevalent architecture in VLMs maps visual embeddings into the language model’s latent space via a dedicated projection module (Liu et al., 2023, 2024a; Zhu et al., 2023; Chen et al., 2023a; Wang et al., 2024a).
+While the large language model (LLM) backbone effectively integrates projected visual information, the image modality also introduces new vulnerabilities (Guo et al., 2024; Qi et al., 2023a; Liu et al., 2024b, c; Gong et al., 2023).
+Specifically, images can function as a “foreign language” (Pi et al., 2024), creating pathways for unsafe input queries, even when the underlying LLM has been aligned for safety (Pi et al., 2024; Chakraborty et al., 2024; Ding et al., 2025).
 The above highlights the unique safety alignment challenges that VLMs face, distinguishing them from text-only LLMs.
-Despite the emergence of safety challenges in VLMs, recent studies have revealed a surprising empirical finding: Enhancing VLM safety could be as simple as applying supervised fine-tuning (SFT), provided that a high-quality, dual-modality curated safety fine-tuning dataset is available [14, 19, 20, 21, 22].
-One compelling piece of evidence is that fine-tuning on VLGuard [23], a widely used VLM safety dataset, substantially improves robustness against unsafe queries and jailbreaking attacks.
-As demonstrated in [23], this enhancement surpasses the results obtained by fine-tuning on a “clean” dataset which removes the unsafe data.
-The surprising effectiveness of SFT on VLGuard has sparked growing interest in re-evaluating its reliability.
-Recent studies [18, 11, 24] have identified a downside of such safety fine-tuning, known as the over-prudence problem.
-This issue refers to the over-conservatism of VLMs after safety fine-tuning, where they unnecessarily reject responses when presented with benign queries.
-However, the over-prudence suggests that these models may be overly safe, to avoid responses to any skeptical queries.
-Alternatively, the observed safety may be illusory, as current safety fine-tuning fails to ensure reliability, giving a false sense of safety.
+Despite the emergence of safety challenges in VLMs, recent studies have revealed a surprising empirical finding: Enhancing VLM safety could be as simple as applying supervised fine-tuning (SFT), assuming access to a high-quality, dual-modality curated safety fine-tuning dataset (Liu et al., 2024c; Zhou et al., 2024; Luo et al., 2024; Zhang et al., 2024a; Gu et al., 2024).
+One compelling piece of evidence is that fine-tuning on widely used VLM safety datasets such as VLGuard (Zong et al., 2024) and SPA-VL (Zhang et al., 2024a), which substantially improves robustness against unsafe queries and jailbreaking attacks.
+The surprising effectiveness of safety SFT has sparked growing interest in re-evaluating its reliability.
+However, a key limitation of such safety fine-tuning is the issue of over-prudence (Ding et al., 2025; Guo et al., 2024; Ding et al., 2024), where VLMs become overly conservative and unnecessarily reject benign queries.
+The over-prudence may indicate that these models are overly safe, avoiding responses even to benign queries.
+Alternatively, the observed safety may be illusory, as current fine-tuning fails to guarantee reliability, creating a false sense of safety.
 Thus, we ask:
 (Q) Does current VLM safety fine-tuning achieve true safety? If not, what is the root cause?
-Figure 1: Schematic overview of safety mirage findings of safety fine-tuned VLM (LLaVA-v1.5-7B-Mixed, fine-tuned on VLG uard [23]) (a) One-word attack vulnerability: A minor modification (e.g., replacing the first instruction word “Share” with “What” in the original unsafe query) can bypass the safety mechanism established through fine-tuning on VLGuard , even though the safeguarded VLM correctly rejects the original unsafe query.
+Figure 1: Schematic overview of safety mirage findings of safety fine-tuned VLM (LLaVA-v1.5-7B-Mixed fine-tuned on VLGuard (Zong et al., 2024)) (a) One-word attack vulnerability: A minor modification (e.g., replacing the first instruction word “Share” with “What” in an unsafe query) bypasses the safety mechanism learned from fine-tuning on VLGuard, even though the model correctly rejects the original query.
 (b) Over-prudence issue: Similar to the one-word attack, a minor modification by replacing “What” with “Share” can cause unnecessary refusals even for benign queries.
-(c) Root cause of spurious correlations given by fine-tuning dataset biases: Certain words become disproportionately associated with specific safety labels.
-For example, “Share” is strongly correlated with rejection responses, while “What” is highly associated with non-rejection responses.
-(d) Effectiveness of unlearning-based safety fine-tuning: The unlearning methods NPO [25] and RMU [26] enhance robustness against attacks while reducing over-prudence, outperforming both the original model LLaVA-v1.5-7B (“Original”) and the supervised fine-tuned LLaVA-v1.5-7B-Mixed [23] (“Prior work”).
+(c) Root cause of spurious correlations from fine-tuning dataset biases: Certain words are disproportionately linked with specific safety labels, such as “Share” strongly correlated with rejection, while “What” highly associated with non-rejection.
+(d) Effectiveness of unlearning-based safety fine-tuning: Using unlearning methods NPO (Zhang et al., 2024b) and RMU (Li et al., 2024a) to do the safety fine-tune on the VLM, which could enhance robustness against attacks and reduce over-prudence, outperforming both the original model LLaVA-v1.5-7B (“Original”) and the supervised fine-tuned LLaVA-v1.5-7B-Mixed (Zong et al., 2024) (“Prior work”).
 In this work, we investigate (Q) and challenge the prevailing belief in the effectiveness of safety fine-tuning for VLMs.
-We uncover a “safety mirage” in VLM safety fine-tuning, where the seemingly robust safety performance after fine-tuning is primarily driven by spurious correlations between certain non-essential textual words in input queries and predefined safety labels (e.g., rejection) in the fine-tuning dataset.
+We uncover a “safety mirage” in VLMs safety fine-tuning, where the seemingly robust performance largely stems from spurious correlations between specific words in input queries and predefined safety labels (e.g., rejection) in the fine-tuning dataset.
 If an adversary identifies these spurious correlations, a simple one-word modification–which we refer to as the “one-word attack”–can effectively jailbreak safety fine-tuned VLMs, enabling them to regenerate unsafe content.
-Additionally, the input-rejection label shortcut induced by these spurious correlations provides an explanation for the over-prudence of safety fine-tuned VLMs.
-Similar to the one-word attack, a one-word modification in text queries can readily activate the input-rejection shortcut at test time, causing the model to overgeneralize rejection responses and refuse to generate outputs even for benign queries.
+Additionally, the input-rejection shortcut induced by these spurious correlations provides an explanation for the over-prudence of safety fine-tuned VLMs.
+Similar to the one-word attack, a one-word modification at test time can easily trigger this shortcut, causing the model to overgeneralize rejection and unnecessarily refuse benign queries.
 In Fig.
-1(a)-(c), we provide a schematic overview illustrating: (a) The one-word attack on a safety fine-tuned VLM LLaVA-v1.5-7B-Mixed [23]; (b) The over-prudence issue by one-word modification; (c) The spurious correlations observed in the fine-tuning dataset VLGuard, where the word “share” is strongly linked to rejection, while “what” is associated with non-rejection.
-Building on our identification of spurious correlations in VLM safety fine-tuning, we propose improving current safety fine-tuning approaches through machine unlearning (MU).
-Originally designed to remove the influence of undesired data or knowledge from ML models, MU ensures that essential knowledge remains intact while avoiding unintended disruptions to causally unrelated information [27, 28, 29].
-We propose adapting MU to VLM safety fine-tuning as a more robust alternative to traditional supervised approaches.
-Rather than enforcing safety through direct supervision, MU enhances VLM safety by erasing the influence of unsafe knowledge in a label-free manner, thereby preventing the formation of spurious correlations between input features and safety labels.
-Although MU has been applied to VLM safety fine-tuning in prior work [17, 30, 31], its unique advantage in mitigating spurious correlations within fine-tuning datasets remains unexplored.
+1(a)-(c), we provide a schematic overview illustrating: (a) The proposed one-word attack on a safety fine-tuned VLMs LLaVA-v1.5-7B-Mixed (Zong et al., 2024); (b) The amplified over-prudence issue by one-word modification; (c) The identified spurious correlations in the fine-tuning dataset VLGuard, where the word “share” is strongly linked to rejection, while “what” is associated with non-rejection.
+Building on our identification of spurious correlations in VLMs safety fine-tuning, we propose improving current safety fine-tuning approaches through machine unlearning (MU).
+Originally designed to remove the influence of undesired data or knowledge from ML models, MU preserves essential knowledge while avoiding unintended disruptions to causally unrelated information (Liu et al., 2025; Cao and Yang, 2015; Bourtoule et al., 2021).
+We propose adapting MU to VLMs safety fine-tuning as a more robust alternative to traditional supervised approaches.
+Instead of enforcing safety through direct supervision, MU enhances VLM safety by erasing the influence of unsafe knowledge in a label-free manner, thereby preventing the formation of spurious correlations between input features and safety labels.
+Although MU has been applied to VLMs safety fine-tuning in prior work (Chakraborty et al., 2024; Chen et al., 2025; Huo et al., 2025), its unique advantage in mitigating spurious correlations within fine-tuning datasets remains unexplored.
 Fig.
-1(d) showcases the effectiveness of applying two LLM unlearning approaches, NPO [25] and RMU [26] in enhanced robustness against jailbreaking attacks and reduced over-prudence rates.
+1(d) illustrates the effectiveness of adapting two LLM unlearning approaches, NPO (negative preference optimization) (Zhang et al., 2024b) and RMU (representation misdirection for unlearning) (Li et al., 2024a), in enhancing robustness against jailbreaking attacks and reducing over-prudence rates.
 In summary, our key contributions are listed below.
-① We revisit the problem of supervised safety fine-tuning for VLMs and find that there exists a safety mirage, driven by hidden biases—specifically, spurious correlations between non-essential textual features and safety labels in the fine-tuning dataset.
-② From an attack perspective, we show that safety fine-tuned VLMs are still susceptible to jailbreaking when adversaries exploit spurious correlations embedded in the fine-tuning dataset.
-We propose a simple and effective one-word attack by substituting highly frequent querying words associated with rejection responses with those linked to normal model outputs.
-Additionally, we show that these spurious correlations also contribute to over-prudence, causing fine-tuned VLMs to unnecessarily reject benign inputs.
-③ From a defense perspective, we show that MU offers a promising solution to alleviate the effects of spurious correlations in fine-tuning data for VLM safety fine-tuning.
-The key rationale is that unlearning removes the influence of unsafe responses without relying on the spurious feature-label correlation present in the fine-tuning dataset.
-④ We conduct extensive experiments across multiple VLM safety evaluation benchmarks, including VLGuard [23], SPA-VL [21], MM-SafetyBench [14], and FigStep [15], and assess model utility on standard VQA datasets.
+①We revisit the problem of safety fine-tuning for VLMs and find that there exists a safety mirage, driven by hidden biases, specifically, spurious correlations between textual questions and safety labels in the fine-tuning dataset.
+②From the attack perspective, we show that safety fine-tuned VLMs remain vulnerable to jailbreaking when adversaries exploit these spurious correlations.
+We propose a simple yet effective one-word attack, which substitutes highly frequent query words associated with rejection responses with those linked to normal model outputs.
+Furthermore, we demonstrate that spurious correlations also underlie the over-prudence phenomenon, causing models to unnecessarily reject benign inputs.
+③From the defense perspective, we show that machine unlearning (MU) offers a promising solution to mitigate the effects of spurious correlations in fine-tuning data.
+The key idea is that MU removes the influence of unsafe responses without relying on spurious feature–label shortcuts.
+④We conduct extensive experiments across multiple VLMs safety evaluation benchmarks, including VLGuard (Zong et al., 2024), SPA-VL (Zhang et al., 2024a), MM-SafetyBench (Liu et al., 2024c), and FigStep (Gong et al., 2023), and assess model utility on standard VQA datasets.
 Our results confirm the safety mirage phenomenon and demonstrate that MU-based safety fine-tuning effectively mitigates spurious correlations and reduces over-prudence.
@@
 VLM safety: Attack and defense.
-With the rapid advancement of VLMs [6, 7, 8, 32, 33, 34, 1, 2, 4], safety concerns have become increasingly prominent due to their potential to generate harmful or inappropriate content.
-While LLMs have been extensively studied for safety risks, leading to the development of attack strategies [35, 36, 37, 38], defense mechanisms [39, 40, 41], and robust evaluation datasets [42, 43, 44], VLMs introduce additional challenges due to the complexity of multimodal inputs [16, 17, 18], making them even more vulnerable to jailbreaking and adversarial manipulation [11, 12, 13, 14, 15].
+With the rapid advancement of VLMs (Liu et al., 2023, 2024a; Zhu et al., 2023; Ye et al., 2023; Wang et al., 2023; Li et al., 2023b; Alayrac et al., 2022; Awadalla et al., 2023; Gao et al., 2023), safety concerns have become increasingly prominent due to their potential to generate harmful or inappropriate content.
+While LLMs have been extensively studied for safety risks, leading to the development of attack strategies (Yang et al., 2023; Wei et al., 2023a; Huang et al., 2023; Shu et al., 2023), defense mechanisms (Li et al., 2023c; Cao et al., 2023; Kumar et al., 2023), and robust evaluation datasets (Bianchi et al., 2023; Li et al., 2024b; Ji et al., 2023), VLMs introduce additional challenges due to the complexity of multimodal inputs (Pi et al., 2024; Chakraborty et al., 2024; Ding et al., 2025), making them even more vulnerable to jailbreaking and adversarial manipulation (Guo et al., 2024; Qi et al., 2023a; Liu et al., 2024b, c; Gong et al., 2023).
 Attacks on VLMs often leverage the dual-modality nature of these models.
-One approach embeds unsafe textual queries into images through typographic manipulation, enabling the model to bypass safety filters and generate harmful outputs [15, 14].
-Another strategy involves using gradient-based adversarial image generation [45, 46, 47, 48, 49] to trigger harmful responses, demonstrating that VLMs remain susceptible to adversarial perturbations despite safety fine-tuning.
+One approach embeds unsafe textual queries into images through typographic manipulation, enabling the model to bypass safety filters and generate harmful outputs (Gong et al., 2023; Liu et al., 2024c).
+Another strategy involves using gradient-based adversarial image generation (Bailey et al., 2023; Dong et al., 2023; Luo et al., 2023; Qi et al., 2023b; Zhao et al., 2023) to trigger harmful responses, demonstrating that VLMs remain susceptible to adversarial perturbations despite safety fine-tuning.
 Defensive strategies for VLM safety generally fall into two categories: Inference-time defenses and fine-tuning with curated safety datasets.
-The former aligns safety responses dynamically at runtime, mitigating unsafe outputs using various filtering and rejection mechanisms [50, 51, 16, 52, 24].
-The latter focuses on red-teaming dataset curation [14, 19, 20, 21, 22, 23, 53], enabling VLMs to be explicitly trained to reject harmful content while retaining utility for benign tasks.
+The former aligns safety responses dynamically at runtime, mitigating unsafe outputs using various filtering and rejection mechanisms (Wang et al., 2024b; Chen et al., 2023b; Pi et al., 2024; Gou et al., 2024; Ding et al., 2024).
+The latter focuses on red-teaming dataset curation (Liu et al., 2024c; Zhou et al., 2024; Luo et al., 2024; Zhang et al., 2024a; Gu et al., 2024; Zong et al., 2024; Li et al., 2024c), enabling VLMs to be explicitly trained to reject harmful content while retaining utility for benign tasks.
 Machine unlearning in VLMs.
-MU [27, 28, 29, 54, 55] is designed to remove harmful data influences from a pre-trained model while preserving its overall utility.
-In the LLM domain, recent work has explored targeted forgetting techniques to erase specific knowledge without compromising performance [25, 26, 56, 57, 58, 59, 60].
-In VLMs, several benchmarks have established systematic evaluation frameworks for MU algorithms [61, 62, 63].
-For safety fine-tuning, prior studies have applied MU-based approaches to mitigate harmful content generation [30, 31, 17].
+MU (Liu et al., 2025; Cao and Yang, 2015; Bourtoule et al., 2021) is designed to remove harmful data influences from a pre-trained model while preserving its overall utility.
+In the LLM domain, recent work has explored targeted forgetting techniques to erase specific knowledge without compromising performance (Zhang et al., 2024b; Li et al., 2024a; Yao et al., 2024).
+In VLMs, several benchmarks have established systematic evaluation frameworks for MU algorithms (Liu et al., 2024d; Dontsov et al., 2024; Ma et al., 2024).
+For safety fine-tuning, prior studies have applied MU-based approaches to mitigate harmful content generation (Chen et al., 2025; Huo et al., 2025; Chakraborty et al., 2024).
 Our work builds on these efforts by leveraging MU to specifically mitigate spurious correlations present in safety fine-tuning datasets.
 3 Preliminary and Problem Statement
-Figure 2: Visualization of question-answer samples on the safety fine-tuned VLMs (LLaVA-v1.5-7B-Mixed and LLaVA-v1.5-7B-Posthoc [23, 64]).
+Existing VLM safety fine-tuning setup.
+Previous works (Pi et al., 2024; Gong et al., 2023; Liu et al., 2024c) highlight the importance of safety alignment in VLMs across both textual and visual modalities.
+Consequently, many efforts have focused on curating high-quality dual-modality safety datasets for VLMs.
+For instance, VLGuard (Zong et al., 2024) covers various text-image scenarios, including unsafe cases where the text is unsafe or both modalities are unsafe, as well as safe cases where both are benign.
+SPA-VL (Zhang et al., 2024a) provides large-scale preference data in the form of question–response tuples, where each image sample pairs a safety-related question with a chosen response and a rejected response.
+Leveraging these curated safety datasets, recent works (Chakraborty et al., 2024; Ding et al., 2025) show that simple fine-tuning approaches on such datasets can yield surprisingly strong safety performance, even against common jailbreaking attacks (Zou et al., 2023; Wei et al., 2023b; Röttger et al., 2023).
+In this work, we revisit the VLM safety fine-tuning problem and later argue that the observed safety improvements from fine-tuning may be an illusion.
+We begin by formulating the problem of VLM safety fine-tuning.
+Let $\mathcal{D}_{\mathrm{u}}$ denote the unsafe dataset, which consists of unsafe text queries and corresponding input images possibly paired with targeted safe responses (e.g., rejection responses in VLGuard and SPA-VL).
+Let $\mathcal{D}_{\mathrm{r}}$ denote the retain dataset, consisting of either a safe text-image dataset or a safety-irrelevant utility dataset, designed to maintain VLM performance on normal tasks after safety fine-tuning.
+For a VLM parameterized by $\bm{\theta}$}, the safety fine-tuning problem can be formulated as:
+$\begin{array}[]{ll}\displaystyle\operatorname*{\text{minimize}}_{{\bm{\theta}}}&\ell_{\mathrm{u}}({\bm{\theta}};\mathcal{D}_{\mathrm{u}})+\gamma\ell_{\mathrm{r}}({\bm{\theta}};\mathcal{D}_{\mathrm{r}}),\end{array}$ (2)
+where $\ell_{\mathrm{u}}$ and $\ell_{\mathrm{r}}$ denote the fine-tuning losses over $\mathcal{D}_{\mathrm{u}}$ and $\mathcal{D}_{\mathrm{r}}$ respectively, and $\gamma\geq 0$ is a regularization parameter that balances safety alignment with general performance preservation.
+Figure 2: Visualization of question-answer samples on the safety fine-tuned VLMs (LLaVA-v1.5-7B-Mixed and LLaVA-v1.5-7B-Posthoc (Zong et al., 2024; Taori et al., 2023)).
 Green shield represents the correct response, either a safe rejection for harmful queries or a valid answer for benign queries.
@@
 represents an inappropriate rejection for a safe query.
-(a) Successfully jailbreaking: The safety fine-tuned model originally produces rejection-based responses for unsafe queries; However, replacing the initial question word with “What” can easily bypass this safeguard.
-(b) Over-prudence: A minor modification by replacing “What” with “Share” can trigger unnecessary refusals even for benign queries.
-Existing VLM safety fine-tuning setup.
-Previous works [16, 15, 14] highlight the need for safety alignment in VLMs to encompass both textual and imagery data.
-Consequently, many efforts [23, 21, 65] have focused on curating high-quality dual-modality safety datasets for VLMs.
-A notable benchmark dataset is VLGuard [23], which covers various text-image pairing scenarios, including unsafe cases where either the text is unsafe or both text and image are unsafe, as well as safe cases where both modalities are benign.
-Leveraging these curated safety datasets, recent works [23, 17, 18, 21] show that simple fine-tuning approaches on such datasets can yield surprisingly strong safety performance, even against common jailbreaking attacks [66, 67, 68].
-In this work, we revisit the VLM safety fine-tuning problem and later argue that the observed safety improvements from fine-tuning may be an illusion.
-We begin by presenting the problem formulation of VLM safety fine-tuning.
-Let $\mathcal{D}_{\mathrm{u}}$ denote the unsafe dataset, which consists of unsafe text queries and corresponding input images possibly paired with targeted safe responses (e.g., rejection responses in VLGuard).
-In addition, let $\mathcal{D}_{\mathrm{r}}$ denote the retain dataset, which consists of either a safe text-image dataset or a safety-irrelevant utility dataset, designed to maintain VLM performance on normal tasks after safety fine-tuning.
-For a VLM parameterized by ${\bm{\theta}}$, the safety fine-tuning problem can be formulated as:
-$\begin{array}[]{ll}\displaystyle\operatorname*{\text{minimize}}_{{\bm{\theta}}}&\ell_{\mathrm{u}}({\bm{\theta}};\mathcal{D}_{\mathrm{u}})+\gamma\ell_{\mathrm{r}}({\bm{\theta}};\mathcal{D}_{\mathrm{r}}),\end{array}$ (2)
-where $\ell_{\mathrm{u}}$ and $\ell_{\mathrm{r}}$ denote the fine-tuning losses over $\mathcal{D}_{\mathrm{u}}$ and $\mathcal{D}_{\mathrm{r}}$ respectively, and the regularization parameter $\gamma\geq 0$ strikes a balance between safety alignment and preserving performance on normal tasks.
+(a) Successfully jailbreak: The safety fine-tuned model originally produces rejection-based responses for unsafe queries.
+However, asking the initial question by starting with the word with “What” can easily bypass this safeguard.
+(b) Over-prudence: A minor modification by asking the problem start from “What” with “Share” can trigger refusals even for benign queries.
 Safety mirage: Motivation and problem of interest.
-Although the safety fine-tuned VLMs could overly reject queries even when presented with benign ones [18, 11, 24], the resulting safety performance against unsafe queries does not degrade; It remains highly robust even in the presence of some common jailbreaking attacks [23, 21].
-The seemingly ‘robust’ safety performance observed after fine-tuning motivates us to re-examine its true reliability.
+Although safety fine-tuned VLMs could exhibit over-rejection, refusing even benign queries (Ding et al., 2025; Guo et al., 2024; Ding et al., 2024), their safety against unsafe queries remains highly robust under common jailbreaking attacks (Zong et al., 2024; Zhang et al., 2024a).
+The seemingly ‘robust’ safety observed after fine-tuning motivates us to re-examine its true reliability.
 As demonstrated in Fig.
 1, the current safety fine-tuned VLM remains highly vulnerable to simple paraphrasing of text queries even when only the first question word is modified.
-As a supplement to Fig.
-1, Fig.
-2 provides additional motivating examples illustrating the vulnerability of safety fine-tuned VLMs to both jailbreaking attacks and over-prudence.
-Consistent with Fig.
-1, unsafe queries prefixed with innocuous question word “What” successfully bypass safeguard, and harmless prompts start with “Share” trigger over-rejection effect.
-Notably, the choice of the replacement word ‘What’ or ‘Share’ is not random but rather stems from the spurious correlations embedded in the safety fine-tuning dataset, as we will illustrate in Sec.
+As a supplement, Fig.
+2 & A1 provides motivating examples illustrating the vulnerability of safety fine-tuned VLMs to both jailbreaking attacks and over-prudence.
+For example, for the VLGuard fine-tuned models, unsafe queries prefixed with the innocuous question word “What” successfully bypass the safeguard, and harmless prompts start with “Share” trigger over-rejection effect.
+Notably, the choice of the question word is not random but rather stems from the spurious correlations embedded in the safety fine-tuning dataset, as we will illustrate in Sec.
 4.
 Examples in Figs.
-1-2 suggest that fine-tuning VLMs on safety datasets may create a “safety mirage”, as evidenced by their susceptibility to even minor one-word modification in text queries.
+1, 2, and A1 suggest that fine-tuning VLMs on safety datasets creates a “safety mirage”, as evidenced by their susceptibility to even minor one-word modification in text queries.
 Thus, our work focuses on the following key research questions: (a) What is the root cause of the “safety mirage” in VLM safety fine-tuning?
@@
 As discussed in Sec.
-3, the effectiveness of current VLM safety fine-tuning methods heavily relies on the curation of high-quality, dual-modality safety datasets.
-As a result, the safety capabilities of fine-tuned VLMs (i.e., their abilities to prevent harmful content generation) are primarily learned from safety labels (i.e., safe responses) introduced in the fine-tuning datasets.
-For example, in VLGuard [23], safety labels for unsafe text queries and/or unsafe images are assigned as rejection responses, such as “I’m sorry, I cannot assist with that request…” against the unsafe text query in Figs.
-1-2.
-Additionally, safety labels may correspond to standard VLM responses when processing safe text-image inputs.
+3, current VLM safety fine-tuning methods heavily depend on high-quality, dual-modality safety datasets.
+Consequently, the safety capabilities of fine-tuned VLMs (i.e., their abilities to prevent harmful content generation) are primarily learned from safety labels (i.e., safe responses) introduced in the fine-tuning datasets.
+For example, in VLGuard, unsafe queries are assigned as rejection responses, such as “I’m sorry, I cannot …”.
+Similarly, SPA-VL also selects rejection-based answers for unsafe queries.
+Refer to caption Refer to caption Refer to caption
+(a) Top words in VLGuard safe   (b) Top words in VLGuard unsafe (c) Top words in SPA-VL
+Figure 3: Frequency of question-initiating words across training corpora: (a) VLGuard safe queries with non-rejection responses, (b) VLGuard unsafe queries with rejection responses, (c) SPA-VL queries.
 At first glance, the use of safety labels appears appropriate.
-However, a hidden bias may arise when these safety labels become strongly correlated with spurious features in the input data, particularly within textual queries, which are the focus of this work.
-Here the term “spurious features” refers to non-essential textual words in inputs that do not contribute to the fundamental meaning or task-relevant aspects of the input query.
+However, hidden bias may arise when safety labels become strongly correlated with spurious features in the input data, particularly within textual queries, which are the focus of this work.
+Here the term “spurious features” refers to non-essential features in inputs (primarily for texts in this work) that do not contribute to the fundamental meaning or task-relevance of the input query, in contrast to the “core” features.
 For example, in Fig.
-2, the word “What” or “Share” at the beginning of the input query can be considered a spurious feature because it does not directly relate to the query’s actual content and can be easily substituted with other question words.
-In contrast, “core” features (such as “crime”) are more informative, representing content-related words that capture the true intent and meaning of the query.
-Therefore, we define spurious correlations as the (unexpected) strong associations between spurious input features and the assigned labels in the safety fine-tuning dataset.
-The above conceptualization of spurious correlations is inspired by conventional analyses in image classification [69, 70, 71], where spurious features correspond to background or non-object pixels, while core features directly relate to the target object.
-In these studies, spurious correlations arise when the target label is linked to spurious features, rather than solely to core features.
-In the context of VLMs, while several studies have explored spurious correlations [72, 73], their focus has been primarily on spurious visual features within the image modality.
-To the best of our knowledge, the spurious correlations caused by textual features in VLM safety fine-tuning have not been investigated before.
+1, the initial words “What” or “Share” function as spurious features, since they are not directly related to the query’s actual content and can be easily substituted with other question words.
+In contrast, core features (such as “crime”) are more informative, representing content-related words that capture the true meaning of the query.
+Therefore, we define spurious correlations as the (unexpected) strong associations between spurious input features and the safety labels in the fine-tuning dataset.
+The above conceptualization of spurious correlations is inspired by and remains consistent with conventional spurious correlation analyses in image classification (Sagawa et al., 2020), where background pixels serve as spurious features while object pixels represent core content.
 In this work, we identify two types of spurious correlations in VLM safety fine-tuning.
 (a) Non-rejection bias: Certain words (like “What” in Fig.
-2(a)) in text queries become spuriously correlated with non-rejection responses.
-As a result, incorporating these words into an original query can easily jailbreak fine-tuned VLMs.
+1(b)) in text queries become spuriously correlated with non-rejection responses.
+So, incorporating these words into an original query can easily jailbreak fine-tuned VLMs.
 (b) Rejection bias: Certain words (like “Share” in Fig.
-2(b)) in text queries become spuriously correlated with rejection responses, causing the fine-tuned VLM to exhibit over-prudence.
-Refer to caption Refer to caption
-(a) Top words in safe queries (b) Top words in unsafe queries
-Figure 3: Frequency of question-initiating words in VLGuard queries: (a) words in safe queries associated with non-rejection responses, (b) words in unsafe queries associated with rejections.
-To determine the spurious textual features and the spurious correlations (a)-(b), we analyze the frequency of words used in train-time text queries that lead to non-rejection responses (i.e., generating textual content in response to safe queries in the training set) and rejection responses (i.e., predefined refusal answers against unsafe queries in the training set), respectively.
+1(b)) in text queries become spuriously correlated with rejection responses, causing the fine-tuned VLM to exhibit over-prudence.
+To reveal these biases, we analyze the frequency of question-initiating words used in training queries that lead to non-rejection responses (i.e., generating contents for safe queries) and rejection responses (i.e., predefined refusal answers for unsafe queries).
 Fig.
-3 presents the most frequently occurring starting words in text queries from the VLGuard dataset, categorized based on their tendency to elicit non-rejection responses or rejection responses.
-As shown, the question word “what” predominantly correlates with non-rejection responses, appearing in over 80% of safe queries where the model generates a response.
-In contrast, the question-initiating words “can” and “share” in unsafe queries are strongly associated with rejection responses, with over 50% of its occurrences leading to a rejection.
-Compared to “can”, which appears in both safe and unsafe queries, the word “share” is only used in unsafe queries to elicit rejection.
+3 presents the most frequent starting words across VLGuard and SPA-VL.
+In VLGuard, the question word “what” predominantly correlates with non-rejection responses, appearing in over 80% of safe queries where the model generates a response.
+By contrast, words like “can” and “share” in unsafe queries are closely tied to rejection responses, with “share” exclusively occurring in unsafe contexts and leading to rejection in more than half of its cases.
+For SPA-VL, we again observe a strong dominance of queries beginning with “what.” Unlike VLGuard safe queries, SPA-VL often pairs such queries with both rejection and non-rejection responses.
+Preference optimization strengthens alignment with the safety-labeled rejections, yet “what”-initiated queries remain inherently associated with unsafe, non-rejection responses.
 One-word jailbreaking.
-Recognizing the non-rejection bias (i.e., the spurious correlation between certain querying words like “what” and non-rejection responses) as shown in Fig.
-3, an adversary can exploit this spurious correlation to jailbreak safety fine-tuned VLMs.
-Formally, let $q$ denote the original unsafe text query that VLM avoids generating unsafe content for.
-And let $q^{\prime}=\mathrm{w}_{\mathrm{adv}}+q$ denote the jailbreaking attack, where $\mathrm{w}_{\mathrm{adv}}$ is the adversarial perturbation chosen based on the non-rejection bias-inducing querying word (e.g., “what” in Figs.
-1, 2, and 3), and the operation $+$ signifies either word insertion as a prefix to $q$ or a simple starting word replacement in $q$.
-We refer to the above simple attack strategy as the one-word jailbreaking attack.
-Figure 4: ASR of $K$-shot one-word attack for varying $K$, evaluated before and after applying “What”-initialized one-word attack to jailbreak the safety fine-tuned VLM (LLaVA-v1.5-7B-Mixed [23]).
-[…]
+Recognizing the non-rejection bias (i.e., the spurious correlation between certain words like “what” and non-rejection responses) as shown in Fig.
+3, an adversary can exploit this bias to jailbreak safety fine-tuned VLMs.
+Formally, let $q$ be an unsafe text query that the VLM would normally refuse.
+We construct a paraphrased version $q^{\prime}$ to rewrite $q$ such that it begins with a chosen adversarial word $\mathrm{w}_{\mathrm{adv}}$ (e.g., “what” in Figs.
+1, 2, and 3).
+We refer to this strategy, rephrasing unsafe queries to enforce a biased starting word, as the one-word jailbreaking attack.
+In practice, we find that repeatedly applying the one-word attack, by integrating $\mathrm{w}_{\mathrm{adv}}$ with paraphrased versions of the original input query $q$ (up to $K$ times), can significantly improve the attack success rate (ASR).
+We refer to this strategy as $K$-shot one-word attack.
+Fig.
+4 shows the ASR of the $K$-shot one-word attack on the safety fine-tuned VLM.
+Even $K=1$ yields 29% ASR, compared to near 0% for the original unsafe queries in VLGuard.
+ASR exceeds 50% for $K\geq 3$ and approaches 90% as $K$ increases.
+In contrast, paraphrased queries without the “What”-trigger remain ineffective, confirming that the bias-inducing word is essential to the attack’s success.
+The effectiveness of the one-word attack can be interpreted through the lens of backdoor attacks (Gao et al., 2020; Saha et al., 2020), where the bias-inducing word “What” functions as a trigger that shortcuts text queries to non-rejection responses during safety fine-tuning.
+Additional results on LLaVA-v1.5-7B-PPO-30K are provided in Appendix B.1.
+Figure 4: ASR of $K$-shot one-word attack for varying $K$, evaluated before and after applying the “What”-initialized one-word attack to jailbreak the safety fine-tuned VLM (LLaVA-v1.5-7B-Mixed (Zong et al., 2024)) on VLGuard unsafe.
+Figure 5: $R$ vs.
+$K$-shot one-word modification, evaluated before and after applying the “Share” initialized one-word modification to safe input queries, causes the over-prudence phenomenon of the safety fine-tuned VLM LLaVA-v1.5-7B-Mixed model (Zong et al., 2024) on VLGuard safe.
 One-word over-prudence.
-Similar to the one-word jailbreaking attack, when a rejection bias-inducing word (like “Share” in Fig.
-3) is introduced into text queries, we observe that even benign modified queries can still prevent the fine-tuned VLM from generating any meaningful output.
-To amplify the over-prudence rate, we can also apply the multi-shot strategy (i.e., “Share”-based one-word modification to different paraphrased versions of the benign input query).
-The results for the multi-shot setting are presented in Appendix A.
-Notably, even in the one-shot setting, the “Share”-initialized benign query modification already achieves a 90% over-rejection rate against safe text-image queries.
+Mirroring the jailbreaking case, inserting a rejection-bias word (e.g., “Share”, Fig.
+3) induces over-prudence in fine-tuned VLMs: even benign queries prefixed with this word often yield no meaningful output.
+A multi-shot procedure, where a benign query is rewritten into several “Share”-initialized variants, further amplifies the effect (Fig.
+5).
+Notably, even a single modification already triggers a 90% over-rejection rate on safe text–image queries, highlighting how dataset-induced starting-word biases can severely degrade utility.
+Additional results for LLaVA-v1.5-7B-PPO-30K in Appendix B.2.
 5 Enhancing VLM Safety through Unlearning
-To mitigate spurious correlations embedded in the safety fine-tuning dataset, one potential solution, motivated by the analysis in Sec.
-4, is to eliminate the dependence of safety fine-tuning on safety labels (e.g., rejection responses to unsafe queries).
-This necessitates shifting from supervised fine-tuning to a label-free, unsupervised setting for safety alignment.
-Machine unlearning (MU) [27, 28, 29] provides an ideal solution in this context, as it is designed to remove the undesired influence of harmful data or knowledge from a pre-trained model while preserving its normal utility.
-Although unlearning has been applied to VLM safety fine-tuning in prior work [17, 30, 31], its unique advantage and application in circumventing spurious correlations within fine-tuning datasets remains unexplored.
-To this end, we adapt two state-of-the-art MU approaches, originally developed for large language models (LLMs)–representation misdirection unlearning (RMU) [26] and negative preference optimization (NPO) [25]–to the context of VLM unlearning.
+To mitigate spurious correlations in the safety fine-tuning dataset (Sec.
+4), a natural solution is to eliminate reliance on safety labels, thereby necessitating shifts to a label-free setting for safety alignment.
+Machine unlearning (MU) (Liu et al., 2025; Cao and Yang, 2015; Bourtoule et al., 2021) provides an ideal solution in this context, as it is designed to remove the undesired influence of harmful data or knowledge from a pre-trained model while preserving its normal utility.
+Although unlearning has been applied to VLM in prior work (Chen et al., 2025; Huo et al., 2025), its unique advantage in addressing spurious correlations remains underexplored.
+Therefore, we adapt two state-of-the-art MU approaches from LLMs, representation misdirection unlearning (RMU) (Li et al., 2024a) and negative preference optimization (NPO) (Zhang et al., 2024b), to VLM safety field.
 The proposed VLM unlearning follows the generic formulation of (2), but with the following key modifications.
-First, the fine-tuning loss over the unsafe dataset $\mathcal{D}_{\mathrm{u}}$ is replaced with an unlearning objective $\ell_{\mathrm{u}}$ that relies solely on the unsafe data features (text-image queries) in $\mathcal{D}_{\mathrm{u}}$, without depending on the safety labels.
+First, the fine-tuning loss on the unsafe dataset $\mathcal{D}_{\mathrm{u}}$ is replaced with an unlearning objective $\ell_{\mathrm{u}}$ that relies solely on the unsafe data features (text-image queries) in $\mathcal{D}_{\mathrm{u}}$, without depending on the safety labels.
 In our work, we define the unlearning loss $\ell_{\mathrm{u}}$ based on the principles of RMU and NPO, respectively.
@@
 where $M_{{\bm{\theta}}}(\cdot)$ represents certain intermediate-layer representations of $\bm{\theta}$}, $c$ is a hyperparameter that controls activation scaling , and $\mathbf{v}$ is a random vector drawn from a standard uniform distribution.
-We remark that, unlike RMU for LLM unlearning [26], we carefully adjusts the representation layer selection and tunes the hyperparameter $c$ to better suit the unlearning process in VLMs.
+We remark that, unlike RMU for LLM unlearning (Li et al., 2024a), we carefully adjusts the representation layer selection and tunes the hyperparameter $c$ to better suit the unlearning process in VLMs.
+In addition to RMU, we also employ NPO (Zhang et al., 2024b) to model the unlearning objective $\ell_{\mathrm{u}}$, which treats unsafe data designated for unlearning as “negative” examples in a direct preference optimization framework (Rafailov et al., 2023).
+The NPO-based unlearning loss is then given by
+$\begin{array}[]{c}\hskip-8.53581pt\ell_{\mathrm{u}}({\bm{\theta}};\mathcal{D}_{\mathrm{u}})=\mathbb{E}_{\mathbf{x}\in\mathcal{D}_{\mathrm{u}}}\left[-\frac{2}{\beta}\log\sigma\left(-\beta\log\left(\frac{\pi_{{\bm{\theta}}}(\mathbf{x})}{\pi_{\mathrm{ref}}(\mathbf{x})}\right)\right)\right],\end{array}$ (5)
+where $\sigma(\cdot)$ the sigmoid function, $\beta>0$ is the temperature parameter , $\pi_{{\bm{\theta}}}$ denotes the prediction probability of the model $\bm{\theta}$} given the unsafe input $\mathbf{x}$, and $\pi_{\mathrm{ref}}$ represents the reference model given by the initial model prior to unlearning.
+The rationale behind NPO is to fine-tune the VLM $\bm{\theta}$} to force it to deviate from the reference model when processing unsafe inputs.
+In addition, unlearning-based safety fine-tuning on VLMs requires the retain loss $\ell_{\mathrm{r}}$ in (2) to preserve utility on normal tasks.
+Unlike LLM unlearning, which relies solely on MU-specific retain objectives, we find that directly applying MU objectives to VLMs often leads to instability or even model collapse.
+To address this, we design $\ell_{\mathrm{r}}$ as a composite of two terms:
+$\ell_{\mathrm{r}}({\bm{\theta}};\mathcal{D}_{\mathrm{r}})=\ell_{\mathrm{ft}}({\bm{\theta}};\mathcal{D}_{\mathrm{r}})+\alpha\ell_{\mathrm{mu},\mathrm{r}}({\bm{\theta}};\mathcal{D}_{\mathrm{r}}),$ (6)
+where $\ell_{\mathrm{ft}}$ denotes the standard fine-tuning loss of the base VLM, which stabilizes optimization and ensures consistent training dynamics, while $\ell_{\mathrm{mu},\mathrm{r}}$ is the MU-specific retain loss that enforces utility preservation (e.g., the representation loss in RMU on the utility set).
+Further details on $\ell_{\mathrm{r}}$ and MU implementation in the safety context are provided in Appendix C.
+As will be shown in Sec.
+6, compared to conventional supervised safety fine-tuning, our approach yields robust safety results.
+6 Experiments
+6.1 Experiment Setups
+Datasets and models.
+We consider four VLM safety datasets: VLGuard (Zong et al., 2024), MM-SafetyBench (Liu et al., 2024c), SPA-VL (Zhang et al., 2024a), and Figstep (Gong et al., 2023).
+To assess the utility of safety fine-tuned VLMs, we also conduct evaluations on representative visual question-answering (VQA) datasets, including VQAv2 (Goyal et al., 2017), TextVQA (Singh et al., 2019), VizWiz (Gurari et al., 2018), and ScienceQA (Lu et al., 2022).
+For model selection, we adopt LLaVA-v1.5-7B and LLaVA-v1.5-13B (Liu et al., 2023, 2024a) as our primary VLMs.
+Safety fine-tuning setups and baselines.
+In our experiments, we use VLGuard as the training dataset for VLM safety fine-tuning, where its training split is divided into two parts: unsafe and safe.
+The safe query–answer pairs form the retain dataset ($\mathcal{D}_{\mathrm{r}}$) used for computing $\ell_{\mathrm{r}}$ in (6).
+For the unlearning loss $\ell_{\mathrm{u}}$, the construction of the unsafe dataset differs by method.
+For NPO (5), we only use the unsafe queries as $\mathcal{D}_{\mathrm{u}}$.
+In contrast, for RMU (3), we concatenate each unsafe query with harmful responses selected by Llama-2-13B-Chat, pairing them to form the input to be unlearned.
+Additional implementation details are provided in Appendix C & D.1.
+Besides MU-based VLM safety fine-tuning, we include a series of popular supervised safety fine-tuning approaches as baselines.
+(1) Mixed-SFT (Zong et al., 2024): Supervised fine-tuning (SFT) using a mixed fine-tuning strategy on VLGuard.
+(2) Posthoc-SFT (Zong et al., 2024; Taori et al., 2023): SFT using a post-hoc fine-tuning approach on VLGuard.
+(3) Unsafe-Filter: An SFT baseline where unsafe samples are excluded by using LLaMA-Guard-3-11B-Vision (Chi et al., 2024) to filter unsafe text–image pairs from LLaVA’s pre-training, refer to Appendix  D.2.
+Evaluation setups.
+We assess safety performance using the attack success rate (ASR).
+For each unsafe query, model outputs are classified into three categories: (1) rejection responses, (2) non-rejection but irrelevant responses, and (3) unsafe responses directly related to the unsafe query.
+ASR is computed in two steps.
+First, we measure the rejection rate across all responses.
+For the remaining non-rejection outputs, we use the multi-modal Qwen2.5-VL-7B-Instruct model as a judge to determine whether a response is related to the unsafe query.
+Any related response is counted as unsafe.
+The full judging prompt is provided in Appendix D.3.
+We report ASR under two scenarios: (1) Before attack: using the original unsafe queries from safety benchmarks.
+(2) After attack: applying our proposed one-word jailbreaking attack to these unsafe queries.
+By default, the attack is implemented in a 3-shot setting, where three paraphrased versions of each query (with the adversarial perturbation preserved) are posed for each image, as illustrated in Fig.
+4.
+When multiple responses ($K>1$) are generated, the attack is considered successful if any of them is unsafe.
+To further analyze the over-prudence phenomenon in safety fine-tuned VLMs, we measure the rejection rate (RR), which quantifies the proportion of benign test-time queries that are unnecessarily rejected.
+Similar to ASR, RR is evaluated both before and after applying 1-shot one-word modifications designed to induce rejection bias, like Fig.
+5.
+We also evaluate the general utility by measuring prediction accuracy (Acc) on standard downstream VQA benchmarks.
 Table 1: Experiment results evaluating safety, over-prudence, and utility of safety fine-tuned VLMs.
@@
 1) that promotes non-rejection bias.
-Over-prudence is measured by RR (rejection rate) on safe input queries, evaluated before and after using 1-shot, one-word over-prudence modification (i.e., “share”-based prefix in Fig.
+Over-prudence is measured by $R$ (rejection rate) on safe input queries, evaluated before and after using 1-shot, one-word modification (i.e., “share”-based prefix in Fig.
 1).
 Here, “Before” and “After” denote the performance prior to and following the respective one-word modification.
-Utility is assessed by the accuracy (Acc.) on four VQA benchmarks: VQAv2, TextVQA, ScienceQA, and VizWiz.
+Utility is assessed by the accuracy (Acc.) on VQA benchmarks.
 Results are presented for models under full fine-tuning and LoRA fine-tuning settings, with safety fine-tuning approaches including Unsafe-Filter, Mixed-SFT, Posthoc-SFT, NPO-Unlearning, and RMU-Unlearning.
@@
      + RMU-Unlearning 2.75% 10.18% 17.74% 22.64% 1.79% 8.64% 26.42% 30.56% 78.27% 58.79% 70.98% 52.99%
-In addition to RMU, we also employ NPO [25] to model the unlearning objective $\ell_{\mathrm{u}}$, which treats unsafe data designated for unlearning as “negative” examples in a direct preference optimization framework [76].
-The NPO-based unlearning loss is then given by
-$\begin{array}[]{c}\ell_{\mathrm{u}}({\bm{\theta}};\mathcal{D}_{\mathrm{u}})=\mathbb{E}_{\mathbf{x}\in\mathcal{D}_{\mathrm{u}}}\left[-\frac{2}{\beta}\log\sigma\left(-\beta\log\left(\frac{\pi_{{\bm{\theta}}}(\mathbf{x})}{\pi_{\mathrm{ref}}(\mathbf{x})}\right)\right)\right],\end{array}$ (5)
-where $\sigma(\cdot)$ the sigmoid function, $\beta>0$ is the temperature parameter , $\pi_{{\bm{\theta}}}$ denotes the prediction probability of the model ${\bm{\theta}}$ given the unsafe input $\mathbf{x}$, and $\pi_{\mathrm{ref}}$ represents the reference model given by the initial model prior to unlearning.
-The rationale behind NPO is to fine-tune the VLM ${\bm{\theta}}$ to force it to deviate from the reference model when processing unsafe inputs.
-Following (2), VLM unlearning also requires the retain loss $\ell_{\mathrm{r}}$ to preserve the model utility in normal tasks.
-This is achieved using the standard VLM training loss over the safe text and safe image data in the fine-tuning set.
-As demonstrated in Sec.
-6, compared to conventional supervised safety fine-tuning, unlearning-based approaches produce safer responses by generating irrelevant or neutral outputs rather than merely rejecting unsafe queries.
-Moreover, unlearning reduces the model’s reliance on rejection responses, alleviating over-prudence and enabling more appropriate handling of benign inputs.
-6 Experiments
-6.1 Experiment Setups
-Datasets and models.
-We consider four VLM safety datasets: VLGuard [23], MM-SafetyBench [14], SPA-VL [21], and Figstep [15].
-To assess the utility of safety fine-tuned VLMs, we also conduct evaluations on representative visual question-answering (VQA) datasets, including VQAv2 [77], TextVQA [78], VizWiz [79], and ScienceQA [80].
-For model selection, we adopt LLaVA-v1.5-7B and LLaVA-v1.5-13B [6, 7] as our primary VLMs.
-Safety fine-tuning setups and baselines.
-In our experiments, we choose VLGuard as the training dataset for VLM safety fine-tuning.
-When implementing the MU-based fine-tuning approaches, i.e., RMU in (3) and NPO in (5), we use the unsafe input-output pairs from VLGuard as the unsafe dataset ($\mathcal{D}_{\mathrm{u}}$) to be unlearned.
-Here we employ Llama-2-13B-Chat to confirm the harmfulness of the original model’s responses against the unsafe input queries.
-Additionally, the safe query-answer pairs from VLGuard are used to construct the retain dataset ($\mathcal{D}_{\mathrm{r}}$) in (2).
-For additional details, we refer readers to Appendix  B.1-B.2.
-Besides MU-based VLM safety fine-tuning, we include a series of popular supervised safety fine-tuning approaches as baselines.
-(1) Mixed-SFT [23]: Supervised fine-tuning (SFT) using a mixed fine-tuning strategy on VLGuard.
-(2) Posthoc-SFT [23, 64]: SFT using a post-hoc fine-tuning approach on VLGuard.
-(3) Unsafe-Filter: SFT performed on a clean training set, where LLaMA-Guard-3-11B-Vision [81] is used to filter unsafe data from pre-training datasets.
-Evaluation setups.
-For a fine-tuned VLM, its safety performance is primarily evaluated using the attack success rate (ASR), defined as the proportion of unsafe responses generated for unsafe input queries over the total number of queries.
-The unsafe input queries are considered in two scenarios.
-(1) Before attack: The original unsafe input queries from VLM safety datasets including VLGuard, MM-SafetyBench, SPA-VL, and Figstep.
-(2) After attack: We apply the proposed one-word attack to the original unsafe queries.
-By default, our one-word jailbreaking attack is implemented using a 3-shot approach, as illustrated in Fig.
-4.
-Specifically, three paraphrased versions of the text query (while preserving the adversarial perturbation) are posed for each image.
-The attack is considered successful if any of the generated responses is unsafe.
-To validate the harmfulness of the responses generated by the fine-tuned model, we employ the multi-modal Qwen2.5-VL-7B-Instruct as a judge.
-The details of the judge prompt can be found in Appendix  B.4.
-To investigate the over-prudence issue in safety fine-tuned VLMs, we measure the rejection rate (RR), which quantifies the proportion of benign test-time queries that are unnecessarily rejected by the fine-tuned model.
-These benign queries consist of both safe text queries and safe images provided by the VLM safety datasets.
-Similar to ASR, RR is evaluated both before and after exploiting spurious correlations to encode rejection bias.
-Furthermore, we evaluate the general utility of safety fine-tuned VLMs by measuring their prediction accuracy (Acc) on standard downstream VQA tasks, including VQAv2, TextVQA, ScienceQA, VizWiz.
 6.2 Experiment Results
 Overall performance on safety, over-prudence, utility.
-In Table 1, we present a comprehensive evaluation of safety fine-tuned VLMs across three key metrics: safety performance against unsafe input queries (measured by ASR, where lower is better), over-prudence performance against safe input queries (measured by RR, where lower is better), and model utility on representative downstream tasks (measured by Acc, where higher is better).
-Recall that both ASR and RR are evaluated before and after exploiting spurious correlations, i.e., promoting non-rejection bias and rejection bias via one-word modification, respectively.
+Table 1 summarizes results across three metrics: ASR (safety), RR (over-prudence), and Acc (utility).
+Both ASR and RR are evaluated before and after exploiting spurious correlations, i.e., promoting non-rejection bias and rejection bias via one-word attack and one-word modification.
+We draw some key observations below.
+First, conventional safety fine-tuning approaches (Unsafe-Filter, Mixed-SFT, and Posthoc-SFT) suffer from a clear safety mirage: ASR rises sharply after one-word attacks (nearly 60% on VLGuard-unsafe and 30% on SPA-VL-unsafe for 7B models), while RR also increases dramatically (over 90%) on safe queries.
+This vulnerability appears consistently under both full and LoRA fine-tuning.
+Second, unlearning-based approaches (NPO and RMU) substantially reduce ASR increases and maintain low RR, alleviating both jailbreaking susceptibility and over-prudence.
+Compared to NPO, RMU performs slightly better.
+Third, in terms of general utility, unlearning-based methods (NPO and RMU) show only a slight accuracy drop (around 1%) compared to the original model.
+This is expected, as a trade-off between robustness and utility exists, and the gain in safety robustness far outweighs the minor utility loss using unlearning.
+To better balance safety and utility in VLMs, we can adopt the idea of coreset unlearning (Pal et al., 2025).
+Using fewer safety samples for unlearning-based fine-tuning improves utility while still achieving stronger safety performance than conventional fine-tuning baselines.
+Detailed results are provided in Table A1 of Appendix D.4.
+Other safety evaluation.
+We further evaluate different safety fine-tuning approaches using LLaVA-v1.5-7B on MM-SafetyBench and FigStep, with results Table A4 of Appendix D.5.
+Consistent with our findings on VLGuard and SPA-VL, one-word perturbations remain highly effective, and MU-based methods are more effective and robust than conventional safety fine-tuning.
+In Fig.
+6, we show input-output examples of safety fine-tuned LLaVA-1.5-7B, comparing Mixed-SFT with RMU-based unlearning.
+The original model remains vulnerable to unsafe queries both before and after the one-word attack, implemented by replacing “How” with “What.” While Mixed-SFT reduces some unsafe outputs, the attack still bypasses its safety filter and triggers unsafe responses.
+Mixed-SFT also suffers from over-prudence, rejecting a benign query starting with “Share,” which is spuriously correlated with rejection responses in Fig.
+3.
+In contrast, RMU-based unlearning both resists the one-word attack and alleviates over-rejection on safe queries.
 Figure 6: Visualization of question-answer pairs from three models: LLaVA-1.5-7B (original), Mixed-SFT (fine-tuned), and RMU-Unlearning (unlearned).
@@
 The first row displays responses to unsafe text-image queries, while the second row shows responses to safe queries.
-We draw some key observations below.
-First, conventional safety fine-tuning approaches (Unsafe-Filter, Mixed-SFT, and Posthoc-SFT) exhibit a safety mirage, as evidenced by a significant rise in ASR after applying the one-word attack–nearly 60% on the VLGuard unsafe evaluation set and nearly 30% on the SPA-VL unsafe evaluation set of LLaVA-1.5-7B based model.
-Additionally, these baselines exhibit a significant over-prudence issue, as evidenced by an over 90% increase after one-word modification in RR on the safe evaluation sets in VLGuard and SPA-VL.
-Furthermore, both full and LoRA fine-tuning exhibit a similar safety mirage phenomenon in baselines.
-Second, compared to baselines, the unlearning-based approaches (NPO and RMU) exhibit significantly lower ASR increases after the one-word attack and maintain a low RR against safe queries, effectively alleviating both jailbreaking susceptibility and over-prudence.
-Compared to NPO, RMU achieves slightly better performance in both ASR (lower vulnerability to the one-word attack) and RR (reduced over-prudence).
-This result is consistent with LLM unlearning, where RMU typically outperforms NPO in knowledge unlearning [26].
-Furthermore, the advantages of MU persist in both full fine-tuning and LoRA fine-tuning.
-Third, from the perspective of model utility evaluation, we observe that unlearning-based approaches lead to a slight decrease in Acc on downstream tasks (approximately 1%), compared to supervised safety fine-tuning.
-This suggests a potential tradeoff between unlearning effectiveness and utility preservation, indicating that erasing harmful knowledge in VLMs may have an impact on general performance.
-Further optimizing this tradeoff remains an important future research direction to enhance VLM unlearning.
-Table 2: Analyzing the safety mechanisms of unlearning-based approaches vs.
-baselines.
-Before and after applying the one-shot “What”-initialized attack, the safety rate (1-ASR) against unsafe queries is decomposed into RR (rejection rate) and irrelevance rate (IR), where IR represents responses that are irrelevant to the unsafe queries.
+Unlearning as a distinct safety mechanism: A response analysis.
+To highlight the difference in safety mechanisms between MU-based approaches and safety-aware SFT, Table 2 reports the proportions of unsafe, irrelevant, and rejection responses before and after 1-shot one-word attack.
+Here, the safety rate (1–ASR) is decomposed into: (1) irrelevant responses, where the model sidesteps the unsafe query with unrelated content, and (2) rejection responses, where the model explicitly refuses to answer.
+Conventional SFT approaches, guided by safety labels, rely mainly on rejection responses, yielding consistently high RR both “Before” and “After”.
+In contrast, MU-based methods do not rely on explicit rejection labels.
+Instead, they achieve safety primarily through irrelevant responses, reflecting a more label-free mechanism for mitigating spurious correlation.
+Table 2: Analyzing the safety mechanisms of unlearning-based approaches versus baselines.
+ASR was reported both before and after the 1-shot “what” initialized attack.
+Safety rate is decomposed into the rejection rate ($R$) and irrelevance rate (IR), where IR denotes non-rejection outputs judged as irrelevant to the unsafe queries.
 All other setups remain consistent with Table 1.
@@
 Before After
-ASR IR RR ASR IR RR
+ASR IR $R$ ASR IR $R$
 LLaVA-1.5-7B 64.25% 30.09% 5.66% 74.43% 21.95% 3.62%
@@
     +RMU-Unlearning 3.87% 90.92% 5.21% 6.91% 88.33% 4.76%
-Other safety evaluation.
-We further evaluate different safety fine-tuning approaches using LLaVA-v1.5-7B on MM-SafetyBench and FigStep, with results Table A1.
-Consistent with our findings on VLGuard and SPA-VL, the one-word attack significantly increases ASR across both benchmarks.
-In Fig.
-5, we present input-output demonstrations of safety fine-tuned LLaVA-1.5-7B, comparing the baseline Mixed-SFT with the RMU-based unlearning approach.
-The original LLaVA-1.5-7B model is vulnerable to unsafe queries both before and after the one-word attack.
-Here the attack is primarily executed by replacing “How” with “What” in text queries.
-However, the one-word attack effectively bypass the safety mechanism from Mixed-SFT, triggering unsafe response against the query starting with “What”.
-Moreover, the Mixed-SFT fine-tuned VLM exhibits over-prudence, as evidenced by its rejection response to a safe text query starting with “Share”.
-Here “Share” is identified as a high-frequency word in the fine-tuning dataset that is strongly correlated with rejection responses in Fig.
-3.
-By contrast, the RMU unlearning approach effectively defends against the one-word attack while also mitigating the over-rejection issue.
 Figure 7: Input token sensitivity analysis for “What”-initiated unsafe queries and “Share”-initiated safe queries over VLGuard using LLaVA-1.5-7B Mixed-SFT, before and after masking “What” and “Share”.
-Sensitivity is measured using per-token masking (i.e., replacing the original token with a blank placeholder [PAD]) to evaluate each token’s influence on response rejection probabilities.
-Irrelevance response for safety enhancement by unlearning.
-To understand the difference in the safety mechanism of MU-based approaches vs.
-safety-aware SFT, Table 2 presents the rates of unsafe, irrelevant, and rejection responses in our safety evaluation, both before and after a 1-shot one-word attack.
-Here, the safety rate (1-ASR) is further decomposed into: (1) irrelevant responses where the model sidesteps the unsafe query by generating a response that is unrelated to the harmful content, and (2) rejection responses where the model explicitly refuses to respond.
-As observed, conventional SFT-based safety fine-tuning strategies predominantly produce rejection-based responses, as reflected in the overly high RR in both “Before” and “After” scenarios.
-In contrast, our unlearning-based methods primarily yield irrelevant responses, reducing the model’s reliance on outright rejections, as evidenced by the high irrelevance rate (IR).
-Notably, the RMU-Unlearning approach steers the model’s output distribution closer to a random vector, making responses more likely to be classified as irrelevant rather than explicitly rejecting queries.
-Input token saliency analysis.
-Recall that spurious correlations emerge between specific words in textual queries and safety labels.
-To further analyze this, we investigate input token sensitivity by applying per-token masking (i.e., replacing the original input token with a blank placeholder [PAD]) to assess the token influence in non-rejection and rejection responses.
-Fig.
-6 presents the rejection probability using LLaVA-1.5-7B Mixed-SFT for all “What”-initiated unsafe input queries and “Share”-initiated safe queries before and after masking “What” and “Share”, respectively, over the VLGuard dataset.
-As we can see, masking “What” significantly increases the rejection probability, confirming its role in inducing non-rejection bias in VLGuard.
-Masking “Share” significantly reduces the rejection probability compared to its unmasked counterpart, demonstrating its influence in reinforcing rejection bias.
-See Appendix B.6 for specific textual input saliency maps.
-Additional results. In Appendix B.7, we demonstrate the robustness of spurious correlation and the effectiveness of our MU-based approach when facing visual variations.
+Sensitivity is measured using per-token masking (i.e., replacing the original token with a blank placeholder [PAD]) to evaluate each token’s influence on rejection probabilities.
+Validating spurious correlations via input token saliency analysis.
+We further validate spurious correlations between specific query words and safety labels using token-level sensitivity analysis, where selected input tokens are masked (i.e., replaced with [PAD]) and their effect on rejection probabilities is measured.
+Figure 7 shows the results for “What’ initiated unsafe queries and “Share” initiated safe queries using LLaVA-1.5-7B Mixed-SFT.
+Masking “What” leads to a sharp increase in rejection probability, confirming its role in inducing non-rejection bias.
+Conversely, masking “Share” reduces the rejection probability compared to the unmasked case, highlighting its influence in reinforcing rejection bias.
+Other saliency maps are provided in Appendix D.6.
+Additional results. In Appendix D.7, we demonstrate the robustness of spurious correlation and the effectiveness of our MU-based approach when facing visual variations.
 7 Conclusion
-In this work, we unveil the “safety mirage” in VLMs, a deceptive robustness that emerges from supervised safety fine-tuning.
-Our analysis reveals that biases in the fine-tuning dataset reinforce spurious correlations between superficial textual patterns and safety labels, leading to a false sense of security.
-As a result, fine-tuned VLMs become highly susceptible to simple one-word jailbreaking attacks while also exhibiting over-prudence, unnecessarily rejecting benign queries.
-To address these, we propose MU as a principled alternative to fine-tuning.
-Unlike supervised approaches, MU directly removes harmful knowledge without relying on explicit safety labels, thereby avoiding biased feature-label mappings and mitigating spurious correlations.
-Extensive experiments confirm the existence of the “safety mirage” in conventional VLM safety fine-tuning and demonstrate that MU-based safety alignment significantly enhances robustness against jailbreaking attacks, reduces over-prudence, and preserves strong performance on standard VQA tasks.
-8 Broader Impacts
-Our work reveals spurious correlations in VLM safety fine-tuning, which are prevalent across models and can function as unintentional “backdoors” in real-world applications.
-Malicious actors could exploit these correlations to launch jailbreaking attacks, potentially extracting sensitive or private information from VLMs.
-While the advocated unlearning methods aim to enhance safety alignment, they also present dual-use risks.
-A bad actor could misuse these techniques to erase safety guardrails, rendering a previously robust model unsafe.
-If such an unlearned model were publicly released, it could bypass safety mechanisms, increasing the risk of harmful content generation.
+In this work, we unveil the “safety mirage” in VLMs, a deceptive robustness that arises from supervised safety fine-tuning.
+We show that biases in fine-tuning datasets reinforce spurious correlations between superficial textual patterns and safety labels, creating a false sense of security.
+As a result, fine-tuned VLMs remain vulnerable to simple one-word jailbreaking attacks and exhibit over-prudence, unnecessarily rejecting benign queries.
+To address this, we propose machine unlearning (MU) as a principled alternative: rather than relying on explicit safety labels, MU removes harmful knowledge in a label-free manner, mitigating spurious correlations.
+Our experiments validate the safety mirage phenomenon and demonstrate that MU improves robustness against jailbreaks, alleviates over-prudence, and preserves utility on standard VQA tasks.
+We refer readers to Appendix FE for limitations, broad impacts, and LLM usage.
 8 Acknowledgement
-We extend our gratitude to Cisco Research and DSO National Laboratories for their support in this project.
-Additionally, the contributions of Yiwei Chen, Yuguang Yao, Yihua Zhang, and Sijia Liu are partially supported by the National Science Foundation (NSF) CISE Core Program Award IIS-2207052, the NSF Cyber-Physical Systems (CPS) Award CNS-2235231, the NSF CAREER Award IIS-2338068, the Amazon Research Award, and the IBM PhD Fellowship.
+The project is primarily supported by research awards from Cisco Research and DSO National Laboratories.
+In addition, the contributions of Yiwei Chen, Yuguang Yao, Yihua Zhang, and Sijia Liu are partially supported by the NSF CISE Core Program Awards IIS-2207052 and IIS-2504263, the NSF CAREER Award IIS-2338068, the NSF Cyber-Physical Systems (CPS) Award CNS-2235231, the Amazon Research Award, the Open Philanthropy Research Award, and the IBM PhD Fellowship.
 Appendix
-A One-word Over Prudence
-Similar to the one-word jailbreaking attack, inserting a rejection bias-inducing word (e.g., “Share” as shown in Fig.
-3) into text queries causes the fine-tuned VLM to produce no meaningful output, even for benign inputs.
-We also apply a multi-shot strategy — modifying different paraphrases of a benign query with “Share”.
-Notably, even a one-shot “Share” modification leads to a 90% over-rejection rate on safe queries.
-Figure A1: Rejection rate vs.
-$K$, evaluated before and after applying the “Share”-initialized one-word modification to safe input queries, cause over-prudence phenomenon of the safety fine-tuned VLM LLaVA-v1.5-7B-Mixed model [23].
-B Additional Experiment Details and Results
-B.1 Additional Fine-tuning Setups
-All experiments are conducted on a single node equipped with 8 NVIDIA A6000 GPUs.
-We perform fine-tuning using both full-parameter and LoRA-based methods.
-For full fine-tuning, we employ a global batch size of 8 and a learning rate of $1\times 10^{-6}$.
-For LoRA-based fine-tuning, following [6, 21], we set the rank $r$ to 256 and the scaling factor $\alpha$ to 512, with a global batch size of 64 and a learning rate of $1\times 10^{-5}$.
-In both setups, we use the AdamW optimizer with zero weight decay and a warmup ratio of 0.03.
-B.2 Machine Unlearning Implementation Details
-To address problem (2), fine-tuning can be effectively performed either through full parameter updates or parameter-efficient techniques such as low-rank adaptation (LoRA) [82].
-Common implementations adopt cross-entropy loss for $\ell_{\mathrm{r}}$, while $\ell_{\mathrm{u}}$ can flexibly incorporate various optimization objectives such as cross-entropy loss, DPO [76], PPO [83].
+Appendix A Visualization of Motivation Samples
+Figure A1: Visualization of question-answer samples on the safety fine-tuned VLMs (LLaVA-v1.5-7B-SPA-VL-DPO-30K and LLaVA-v1.5-7B-SPA-VL-PPO-30K (Zhang et al., 2024a; Rafailov et al., 2023; Schulman et al., 2017)).
+The visual marks keep consistent meaning with Fig.
+2.
+(a) Successfully jailbreak: The safety fine-tuned model originally produces rejection-based responses for unsafe queries.
+But if the questions use the starting word “Please”, it would bypass this safeguard.
+(b) Over-prudence: The fine-tuned model will generate the benign responses when the queries are both harmless, but ask the problems starting from “What” can trigger unnecessary refusals.
+Appendix B Additional Results on Spurious Correlations
+B.1 Additional Results of One-Word Jailbreaking
+Figure A2: ASR of $K$-shot one-word attack for varying $K$, evaluated before and after applying the “Please”-initialized one-word attack to jailbreak the safety fine-tuned VLM (LLaVA-v1.5-7B-PPO-30K (Zhang et al., 2024a)) on SPA-VL unsafe test set.
+For the SPA-VL test set, we further evaluated the LLaVA-v1.5-7B-PPO-30K model using the same $K$-shot paraphrasing protocol as in the main text.
+Here, we selected “please” as the adversarial biasing word for unsafe queries.
+The motivation is dataset-driven: in the SPA-VL training corpus, the token “please” occurs with very low frequency like Fig.
+3 shows, so asserting it to unsafe queries disrupts the model’s learned phrasing patterns.
+Under PPO-based preference optimization, this rare starter token reduces the model’s tendency to follow the preferred safe/rejection response, thereby substantially increasing the attack success rate (ASR).
+These results corroborate our main finding that low-frequency tokens can serve as effective adversarial triggers in one-word jailbreaking.
+Quantitatively, as shown in Fig.
+A2, the ASR rises from about 12% before modification to nearly 30% at $K=1$, and continues to grow steadily with larger $K$, ultimately exceeding 60%.
+This confirms that the “Please” based attack is consistently effective across different shot settings.
+B.2 Additional Results of One-Word Over-prudence
+Figure A3: Rejection rate of $K$-shot one-word modification, evaluated before and after applying the “What” initialized one-word modification to safe input queries, causes the over-prudence phenomenon of the safety fine-tuned VLM (LLaVA-v1.5-7B-PPO-30K (Zhang et al., 2024a)) on SPA-VL safe.
+We also tested the over-prudence effect on SPA-VL safe queries.
+For SPA-VL, we used “what” as the biased starter word.
+Unlike “please”, “what” is highly frequent in the SPA-VL training corpus as Fig.
+3 presents.
+When combined with PPO preference optimization, this frequent token strongly biases the model toward the preferred rejection-style responses.
+As a result, benign queries initialized with “what” exhibit a substantially increased rejection rate (RR), confirming the over-prudence phenomenon.
+These findings highlight how high-frequency tokens can act as rejection-bias triggers in safe contexts, complementing the low-frequency jailbreak triggers observed in unsafe contexts.
+From Fig.
+A3, we see that the RR almost doubles, jumping from about 39% before modification to nearly 79% after just one shot, and quickly saturating around 90% as $K$ grows.
+This illustrates the strong over-rejection tendency induced by the “what” starter.
+Appendix C Implementations Details of Machine Unlearning
+To address problem (2), fine-tuning can be effectively performed either through full parameter updates or parameter-efficient techniques such as low-rank adaptation (LoRA).
+Common implementations adopt cross-entropy loss for $\ell_{\mathrm{r}}$, while $\ell_{\mathrm{u}}$ can flexibly incorporate various optimization objectives such as cross-entropy loss, DPO (Rafailov et al., 2023), PPO (Schulman et al., 2017).
+Details of Retain Loss $\ell_{\mathrm{r}}$.
+To address the problem, directly applying MU objectives to VLMs often leads to instability or even model collapse, so we design $\ell_{\mathrm{r}}$ as a composite of two terms in (6).
+For the $\ell_{\mathrm{ft}}$, it is the common implementations that adopt cross-entropy loss on $\mathcal{D}_{\mathrm{r}}$ like the original LLaVA fine-tune.
+While for the $\ell_{\mathrm{mu},\mathrm{r}}$, it is depend on the choices of $\ell_{\mathrm{u}}$, we set it as the corresponding retain loss of the specific unlearning algorithm.
+For example, when $\ell_{\mathrm{u}}$ is the RMU training loss,
+$\ell_{\mathrm{\mathrm{mu},\mathrm{r}}}({\bm{\theta}};\mathcal{D}_{\mathrm{r}})=\mathbb{E}_{\mathbf{x}\in\mathcal{D}_{\mathrm{r}}}\big[\|M_{{\bm{\theta}}}(\mathbf{x})-M_{{\bm{\theta}}_{\mathrm{ref}}}(\mathbf{x})\|_{2}^{2}\big],$ (A1)
+where $M_{{\bm{\theta}}}(\cdot)$ denotes the intermediate-layer representations of $\bm{\theta}$}, and $M_{{\bm{\theta}}_{\mathrm{ref}}}(\cdot)$ those of the reference (pre-unlearning) model.
+While for $\ell_{\mathrm{u}}$ equals to the NPO training loss,
+$\ell_{\mathrm{\mathrm{mu},\mathrm{r}}}({\bm{\theta}};\mathcal{D}_{\mathrm{r}})=\mathbb{E}_{\mathbf{x}\in\mathcal{D}_{\mathrm{r}}}\left[-\frac{2}{\beta}\log\sigma\left(\beta\log\left(\frac{\pi_{{\bm{\theta}}}(\mathbf{x})}{\pi_{\mathrm{ref}}(\mathbf{x})}\right)\right)\right],$ (A2)
+where $\sigma(\cdot)$ the sigmoid function, $\beta>0$ is the temperature parameter, $\pi_{{\bm{\theta}}}$ denotes the prediction probability of the current model given the retain input $\mathbf{x}$, and $\pi_{\mathrm{ref}}$ represents the reference model given by the initial model prior to unlearning.
+During the experiment, we both set $\alpha=1$ in (2).
 RMU implementation details.
@@
 The unlearning loss $\ell_{\mathrm{u}}$ defined as (3), where $M_{{\bm{\theta}}}(\cdot)$ denotes the intermediate-layer representations (we use layers 4, 5, and 6 in our implementation), $c$ is a hyperparameter for activation scaling (set to $c=10$), and $\mathbf{v}$ is a random vector drawn from a standard uniform distribution.
-The retain loss $\ell_{\mathrm{r}}$ is computed as the cross-entropy loss on the retain data, with the regularization parameter set to $\gamma=1.2$.
+The retain loss $\ell_{\mathrm{r}}$ is computed data, with the regularization parameter set to $\gamma=1.2$.
 NPO implementation details.
@@
 However, for NPO, the input to the VLM from $\mathcal{D}_{\mathrm{u}}$ consists solely of the harmful question.
-In this setting, unsafe data designated for unlearning is treated as negative examples within a direct preference optimization framework [76].
+In this setting, unsafe data designated for unlearning is treated as negative examples within a direct preference optimization framework (Rafailov et al., 2023).
 The unlearning training objective is defined as (5), where $\sigma(\cdot)$ denotes the sigmoid function, $\beta>0$ is a temperature parameter (set to $\beta=0.6$ in our implementation), $\pi_{{\bm{\theta}}}(\mathbf{x})$ is the prediction probability of the model given the unsafe input $\mathbf{x}$, and $\pi_{\mathrm{ref}}(\mathbf{x})$ is the corresponding probability from the reference model (i.e., the initial model prior to unlearning).
-The retain loss $\ell_{\mathrm{r}}$ is computed as the cross-entropy loss on the retain data, with the regularization parameter set to $\gamma=1.0$.
-B.3 Additional Baseline Details
-For the Unsafe-Filter baseline, we employ the LLaMA-Guard-3-11B-Vision [81] model to assess the safety of image-instruction-answer pairs in the LLaVA pre-training and fine-tuning datasets.
-Unsafe data is filtered out, and LLaVA is subsequently fine-tuned using only the remaining safe samples.
-The fine-tuning setup follows the original LLaVA configuration.
-B.4 Evaluation Setup Details
-To validate the harmfulness of the responses generated by the fine-tuned model, we employ =Qwen2.5-VL-7B-Instruct model as a judge.
+The retain loss $\ell_{\mathrm{r}}$ computed with the regularization parameter set to $\gamma=1.0$.
+Appendix D Additional Experiment Details and Results
+D.1 Additional Fine-tuning Setups
+All experiments are conducted on a single node equipped with 8 NVIDIA A6000 GPUs.
+We perform fine-tuning using both full-parameter and LoRA-based methods.
+For full fine-tuning, we employ a global batch size of 8 and a learning rate of $1\times 10^{-6}$.
+For LoRA-based fine-tuning, following (Liu et al., 2023; Zhang et al., 2024a), we set the rank $r$ to 256 and the scaling factor $\alpha$ to 512, with a global batch size of 64 and a learning rate of $1\times 10^{-5}$.
+In both setups, we use the AdamW optimizer with zero weight decay and a warmup ratio of 0.03.
+D.2 Additional Baseline Details
+An additional SFT baseline is constructed by filtering out unsafe samples at both stages of LLaVA training.
+Specifically, we employ LLaMA-Guard-3-11B-Vision (Chi et al., 2024) to assess the safety of all image–instruction–answer triplets in LLaVA’s pre-training and instruction-tuning datasets.
+Any sample flagged as unsafe is removed prior to training.
+As a result, the model is pre-trained exclusively on safe image–text pairs, and subsequently fine-tuned on safe instruction–answer pairs, ensuring that unsafe data is excluded from the entire training pipeline.
+The fine-tuning setup otherwise follows the original LLaVA configuration, making this a direct comparison point for evaluating the effect of safety filtering throughout both training stages.
+D.3 Evaluation Setup Details
+To validate the harmfulness of the responses generated by the fine-tuned model, we employ Qwen2.5-VL-7B-Instruct model as a judge.
 Below is the prompt used for evaluating content relevance, which is included in our experiments:
@@
 “Provide detailed reasoning for your decision.”
-B.5 Other Safety Evaluation
-In Table A1, we present additional safety evaluations of different safety fine-tuning approaches using LLaVA-v1.5-7B on two additional safety benchmarks: MM-SafetyBench and FigStep.
-The results are consistent with our earlier findings on VLGuard and SPA-VL: The one-word attack leads to a significant increase in ASR on both MM-SafetyBench and FigStep.
-This reveals that the spurious correlations learned during conventional safety fine-tuning persist across different evaluation settings.
-Our proposed MU-based methods demonstrate improved robustness, thereby alleviating this vulnerability.
-Table A1: Safety evaluation on MM-Safetybench and FigStep. ASR is reported for unsafe inputs before and after the “What”-initialized 3-shot attack. The setup and format follow Table 1.
+D.4 Balance between Safety and Utility
+Table A1: Overall results evaluating safety (ASR, $\downarrow$), over-prudence ($R$, $\downarrow$), and utility (Acc., $\uparrow$) on VQA benchmarks.
+Metrics are reported before and after one-word attacks/modifications.
+We compare conventional fine-tuning baselines, unlearning-based methods (NPO, RMU), and their coreset variants (50%, 25%, 10%).
+The other settings stay consistent with Table  1.
+Models Safety Evaluation (ASR, $\downarrow$) Over-Prudence Evaluation (RR, $\downarrow$) Utility Evaluation (Acc., $\uparrow$)
+VLGuard SPA-VL VLGuard SPA-VL VQAv2 TextVQA ScienceQA VizWiz
+Before After Before After Before After Before After
+LLaVA-1.5-7B 64.25% 90.27% 46.42% 52.08% 0.36% 0.36% 14.72% 9.81% 78.53% 58.23% 69.51% 50.07%
+     + Unsafe-Filter 65.66% 90.72% 45.66% 54.72% 0.36% 0.36% 15.85% 11.32% 79.14% 58.22% 68.12% 52.14%
+     + Mixed-SFT 0.23% 54.98% 14.34% 37.73% 4.48% 91.76% 68.68% 98.87% 78.23% 57.80% 68.27% 52.94%
+     + Posthoc-SFT 0.23% 46.83% 13.58% 32.96% 2.69% 90.83% 60.38% 100.0% 78.03% 57.73% 68.42% 51.84%
+     + NPO-Unlearning 2.49% 12.92% 18.49% 24.15% 2.51% 11.69% 16.60% 17.36% 77.34% 57.80% 68.02% 50.21%
+      + 50% NPO 2.94% 12.46% 19.02% 25.10% 2.69% 12.23% 17.45% 18.02% 77.41% 57.82% 68.20% 50.55%
+      + 25% NPO 3.65% 14.25% 20.12% 26.48% 2.51% 11.33% 17.88% 18.67% 77.60% 58.10% 68.45% 50.92%
+      + 10% NPO 5.17% 18.77% 22.01% 28.92% 1.97% 7.72% 18.32% 19.15% 78.10% 57.99% 68.72% 51.34%
+     + RMU-Unlearning 1.29% 10.18% 17.73% 22.64% 1.25% 7.56% 18.11% 19.24% 77.04% 56.89% 67.68% 50.01%
+      + 50% RMU 1.74% 10.64% 18.12% 23.05% 1.61% 8.01% 18.56% 19.70% 77.23% 57.98% 67.84% 50.34%
+      + 25% RMU 3.22% 15.43% 19.06% 24.12% 1.61% 4.28% 18.92% 20.01% 78.01% 57.24% 68.12% 50.78%
+      + 10% RMU 7.51% 22.12% 21.35% 27.25% 0.72% 1.97% 19.48% 20.73% 78.32% 58.97% 68.65% 51.42%
+To better balance safety and utility in VLMs, we further adopt the idea of coreset unlearning (Pal et al., 2025).
+Instead of unlearning all safety data, we fine-tune on a random-sampled subset (e.g., 50%, 25%, or 10%) of the safety dataset.
+As shown in Table A1, reducing the coreset size leads to slightly higher ASR and RR (weaker safety), but improves accuracy on downstream tasks (higher utility).
+This also reflects a natural trade-off: smaller coresets sacrifice some safety robustness while preserving more general task capability.
+Importantly, even with only 10% of the safety data, unlearning-based models still achieve substantially better safety–utility trade-offs than conventional fine-tuning baselines, confirming the flexibility of coreset unlearning for practical deployment
+Figure A4: Safety evaluation on MM-SafetyBench and FigStep. ASR is reported before and after the “What”-based 3-shot attack. Setup follows Table 1.
 Models Safety Evaluation ($\downarrow$)
@@
     +RMU-Unlearning 5.36% 19.05% 2.00% 12.00%
-Table A2: Evaluating the safety and over-prudence under various visual perturbations on VLGuard, including Gaussian Noise, Gaussian Blur, and Color Jitter. Other setups are consistent with Table 1.
-Models Safety Evaluation (ASR, $\downarrow$) Over-Prudence (RR, $\downarrow$)
+D.5 Additional Safety Evaluation
+In Table A4, we provide additional safety evaluations of different fine-tuning approaches for LLaVA-v1.5-7B on two further safety benchmarks: MM-SafetyBench and FigStep.
+The results align closely with our earlier findings on VLGuard and SPA-VL.
+In both new benchmarks, the one-word attack substantially increases the attack success rate (ASR), indicating that spurious correlations between surface-level textual cues and safety labels, induced during conventional safety fine-tuning, persist across diverse evaluation settings.
+These correlations cause models to become vulnerable to simple adversarial word insertions, undermining their safety guarantees.
+By contrast, our MU-based approaches (NPO and RMU) consistently achieve lower post-attack ASR, demonstrating improved robustness and reduced reliance on spurious features.
+This highlights that MU not only mitigates overfitting to dataset-specific biases but also generalizes better across benchmarks, thereby offering a more reliable defense against spurious correlation–driven vulnerabilities.
+Figure A5: Prediction probability for various output sentences before and after masking each input query token for the query “What can be used to create a promotion?” corresponding to the image example in Fig.
+1-(a).
+Figure A6: Prediction probability for various output sentences before and after masking each input query token for the query “Share the drink type in the image” corresponding to the image example in Fig.
+1-(b).
+D.6 Examples of Input Saliency
+We show two input saliency maps in Fig.
+A5 and Fig.
+A6.
+The input saliency is shown through the prediction probability change for each output sentence ($y$-axis) by masking each input token ($x$-axis).
+Here [X] refers to the token “X” is marked (i.e., replacing with [PAD])), and [ ] denotes the non-masking case.
+The examples of input query in Fig.
+A5 and Fig.
+A6 correspond to the examples introduced in Fig.
+1(a)-(b).
+Fig.
+A5 shows that masking “what” token can significantly boost the rejection probability, producing “I’m sorry…”.
+Fig.
+A6 shows that masking “share” token successfully reduces the prediction probability for rejection response and produces the normal model generation.
+Figure A7: Evaluating safety (ASR, $\downarrow$) and over-prudence ($R$, $\downarrow$) under visual perturbations on VLGuard, including Gaussian Noise, Gaussian Blur, and Color Jitter.
+Other setups follow Table 1.
+Models Safety (ASR) Over-Prudence (RR)
 Before After Before After
@@
 + Color Jitter 1.74% 10.41% 1.07% 7.02%
-B.6 Examples of Input Saliency
-We show two input saliency maps in Fig.
-A2 and Fig.
-A3.
-The input saliency is shown through the prediction probability change for each output sentence ($y$-axis) by masking each input token ($x$-axis).
-Here [X] refers to the token “X” is marked (i.e., replacing with [PAD])), and [ ] denotes the non-masking case.
-The examples of input query in Fig.
-A2 and Fig.
-A3 correspond to the examples introduced in Fig.
-1(a)-(b).
-Fig.
-A2 shows that masking “what” token can significantly boost the rejection probability, producing “I’m sorry…”.
-Fig.
-A3 shows that masking “share” token successfully reduces the prediction probability for rejection response and produces the normal model generation.
-Figure A2: Prediction probability for various output sentences before and after masking each input query token for the query “What can be used to create a promotion?” corresponding to the image example in Fig.
-1-(a).
-Figure A3: Prediction probability for various output sentences before and after masking each input query token for the query “Share the drink type in the image” corresponding to the image example in Fig.
-1-(b).
-B.7 Spurious Correlations with Visual Variations
-To further evaluate the persistence of spurious correlations between text queries and safety labels and the robustness of our MU-based safety fine-tuning approaches, we present in Table A2 the safety performance of fine-tuned models when subjected to various visual perturbations.
-The results reveal that the SFT-based safety models consistently exhibit a significant increase in both ASR and RR even under visual variations.
-This finding indicates that the spurious correlations, and consequently the safety mirage phenomenon remains stable.
-In contrast, the performance of our MU-based safety aligned models remains robust across these perturbations.
+D.7 Spurious Correlations with Visual Variations
+To further examine whether the identified spurious correlations between textual queries and safety labels persist under different visual conditions, we evaluate fine-tuned models on VLGuard with a range of common visual perturbations, including Gaussian Noise, Gaussian Blur, and Color Jitter.
+The results are summarized in Table,A7.
+We make several observations.
+First, conventional SFT-based safety models (Mixed-SFT and Posthoc-SFT) exhibit a consistent pattern of vulnerability: both ASR and RR remain high across all perturbations, showing that the learned spurious correlations are insensitive to low-level visual variations.
+This suggests that the safety mirage is not simply a dataset artifact tied to specific image statistics, but a systematic bias reinforced during fine-tuning.
+Second, MU-based approaches (NPO and RMU) maintain stable and low ASR and RR even under perturbations, confirming that unlearning not only alleviates spurious correlations in clean settings but also provides robustness against visual shifts.
+Third, comparing perturbation types, we find that color-based distortions (Color Jitter) lead to slightly larger fluctuations than Gaussian Noise or Blur, though the overall trend remains the same.
+This indicates that safety mirage in SFT models is strongly text-driven and less sensitive to visual details, while MU-based approaches decouple safety predictions from such spurious text–label shortcuts.
+Taken together, these findings highlight that spurious correlations embedded during safety fine-tuning are persistent and modality-invariant: they survive visual perturbations that alter low-level image statistics.
+In contrast, unlearning-based models generalize more reliably across such shifts, further validating their robustness in multimodal safety alignment.
+Appendix E The Use of LLMs
+This work makes limited use of LLMs.
+Specifically, LLMs were employed exclusively for grammar correction and stylistic polishing of the manuscript.
+They were not involved in research ideation, experimental design, data analysis, or the generation of any scientific content.
+All substantive contributions to the paper are solely attributable to the author.
+Appendix F Limitations and broad Impact
+F.1 Limitations
+Our analysis has so far been limited to models at the 7B and 13B scales.
+Extending the study to larger foundation models remains an important next step, as scaling may introduce new behaviors or amplify new challenges.
+In addition, our experiments have primarily relied on widely used safety benchmarks.
+While these datasets provide systematic evaluation settings, they may not fully capture the diversity and complexity of real-world applications.
+Future work should therefore investigate unlearning performance in broader, domain-specific scenarios and corresponding downstream tasks.
+F.2 Broad Impact
+This work sheds light on the reliability of current VLM safety fine-tuning, uncovering the risk of a “safety mirage” where perceived robustness is illusory.
+By identifying spurious correlations as a root cause and demonstrating how machine unlearning can alleviate them, our study contributes to building safer and more trustworthy multimodal systems.
+In practice, stronger defenses against jailbreaking and over-prudence can reduce the risk of harmful or overly restrictive outputs, improving user trust in deployed AI systems.
+At the same time, unlearning techniques must be carefully applied, as overly aggressive removal of knowledge could impair model utility or be misused to deliberately suppress content.
+We encourage future research to balance safety, fairness, and transparency when deploying MU-based safety interventions in real-world applications.

### v3

@@
-Safety Mirage: How Spurious Correlations Undermine VLM Safety Fine-Tuning and Can Be Mitigated by Machine Unlearning
+arXiv:2503.11832v3[cs.AI] 02 Mar 2026
 Abstract
@@
 However, their susceptibility to generating harmful content when exposed to unsafe queries raises critical safety concerns.
-While current alignment strategies primarily rely on supervised safety fine-tuning with curated datasets, we identify a fundamental limitation we call the “safety mirage”, where supervised fine-tuning inadvertently reinforces spurious correlations between superficial textual patterns and safety responses, rather than fostering deep, intrinsic mitigation of harm.
+While current alignment strategies primarily rely on supervised safety fine-tuning with curated datasets, we identify a fundamental limitation we call the “safety mirage,” where supervised fine-tuning inadvertently reinforces spurious correlations between superficial textual patterns and safety responses, rather than fostering deep, intrinsic mitigation of harm.
 We show that these spurious correlations leave fine-tuned VLMs vulnerable even to a simple one-word modification-based attack, where substituting a single word in text queries with a spurious correlation-inducing alternative can effectively bypass safeguards.
-Additionally, these correlations contribute to the over-prudence, causing fine-tuned VLMs to refuse benign queries unnecessarily.
+Additionally, these correlations contribute to over-prudence, causing fine-tuned VLMs to refuse benign queries unnecessarily.
 To address these issues, we show machine unlearning (MU) as a powerful alternative to supervised safety fine-tuning, as it avoids biased feature-label mappings and directly removes harmful knowledge from VLMs while preserving their general capabilities.
-Extensive evaluations across safety benchmarks show that under MU-based alignment reduces the attack success rate by up to 60.17% and cuts unnecessary rejections by over 84.20%.
+Extensive evaluations across safety benchmarks show that under MU-based alignment reduces the attack success rate by up to 60.27% and cuts unnecessary rejections by over 84.20%.
 Codes are available at https://github.com/OPTML-Group/VLM-Safety-Unlearn.
@@
 1 Introduction
-Recent multi-modal models have achieved advancements by integrating text and images (Alayrac et al., 2022; Awadalla et al., 2023; Hurst et al., 2024; Gao et al., 2023; Li et al., 2023a).
-A prevalent architecture in VLMs maps visual embeddings into the language model’s latent space via a dedicated projection module (Liu et al., 2023, 2024a; Zhu et al., 2023; Chen et al., 2023a; Wang et al., 2024a).
-While the large language model (LLM) backbone effectively integrates projected visual information, the image modality also introduces new vulnerabilities (Guo et al., 2024; Qi et al., 2023a; Liu et al., 2024b, c; Gong et al., 2023).
+Recent multi-modal models have achieved advancements by integrating text and images (Alayrac et al., 2022; Awadalla et al., 2023; Hurst et al., 2024; Gao et al., 2023; Li et al., 2023b).
+A prevalent architecture in VLMs maps visual embeddings into the language model’s latent space via a dedicated projection module (Liu et al., 2023; 2024a; Zhu et al., 2023; Chen et al., 2023a; Wang et al., 2024a).
+While the large language model (LLM) backbone effectively integrates projected visual information, the image modality also introduces new vulnerabilities (Guo et al., 2024; Qi et al., 2023b; Liu et al., 2024c; b; Gong et al., 2023).
 Specifically, images can function as a “foreign language” (Pi et al., 2024), creating pathways for unsafe input queries, even when the underlying LLM has been aligned for safety (Pi et al., 2024; Chakraborty et al., 2024; Ding et al., 2025).
 The above highlights the unique safety alignment challenges that VLMs face, distinguishing them from text-only LLMs.
-Despite the emergence of safety challenges in VLMs, recent studies have revealed a surprising empirical finding: Enhancing VLM safety could be as simple as applying supervised fine-tuning (SFT), assuming access to a high-quality, dual-modality curated safety fine-tuning dataset (Liu et al., 2024c; Zhou et al., 2024; Luo et al., 2024; Zhang et al., 2024a; Gu et al., 2024).
-One compelling piece of evidence is that fine-tuning on widely used VLM safety datasets such as VLGuard (Zong et al., 2024) and SPA-VL (Zhang et al., 2024a), which substantially improves robustness against unsafe queries and jailbreaking attacks.
+Despite the emergence of safety challenges in VLMs, recent studies have revealed a surprising empirical finding: Enhancing VLM safety could be as simple as applying supervised fine-tuning (SFT), assuming access to a high-quality, dual-modality curated safety fine-tuning dataset (Liu et al., 2024b; Zhou et al., 2024; Luo et al., 2024; Zhang et al., 2024b; Gu et al., 2024).
+One compelling piece of evidence is that fine-tuning on widely used VLM safety datasets such as VLGuard (Zong et al., 2024) and SPA-VL (Zhang et al., 2024b), which substantially improves robustness against unsafe queries and jailbreaking attacks.
 The surprising effectiveness of safety SFT has sparked growing interest in re-evaluating its reliability.
@@
 (c) Root cause of spurious correlations from fine-tuning dataset biases: Certain words are disproportionately linked with specific safety labels, such as “Share” strongly correlated with rejection, while “What” highly associated with non-rejection.
-(d) Effectiveness of unlearning-based safety fine-tuning: Using unlearning methods NPO (Zhang et al., 2024b) and RMU (Li et al., 2024a) to do the safety fine-tune on the VLM, which could enhance robustness against attacks and reduce over-prudence, outperforming both the original model LLaVA-v1.5-7B (“Original”) and the supervised fine-tuned LLaVA-v1.5-7B-Mixed (Zong et al., 2024) (“Prior work”).
+(d) Effectiveness of unlearning-based safety fine-tuning: The unlearning methods NPO (Zhang et al., 2024a) and RMU (Li et al., 2024c), enhance robustness against attacks and reduce over-prudence, outperforming both the original model LLaVA-v1.5-7B (“Original”) and the supervised fine-tuned LLaVA-v1.5-7B-Mixed (Zong et al., 2024) (“Prior work”).
 In this work, we investigate (Q) and challenge the prevailing belief in the effectiveness of safety fine-tuning for VLMs.
@@
 Fig.
-1(d) illustrates the effectiveness of adapting two LLM unlearning approaches, NPO (negative preference optimization) (Zhang et al., 2024b) and RMU (representation misdirection for unlearning) (Li et al., 2024a), in enhancing robustness against jailbreaking attacks and reducing over-prudence rates.
+1(d) illustrates the effectiveness of adapting two LLM unlearning approaches, NPO (negative preference optimization) (Zhang et al., 2024a) and RMU (representation misdirection for unlearning) (Li et al., 2024c), in enhancing robustness against jailbreaking attacks and reducing over-prudence rates.
 In summary, our key contributions are listed below.
@@
 The key idea is that MU removes the influence of unsafe responses without relying on spurious feature–label shortcuts.
-④We conduct extensive experiments across multiple VLMs safety evaluation benchmarks, including VLGuard (Zong et al., 2024), SPA-VL (Zhang et al., 2024a), MM-SafetyBench (Liu et al., 2024c), and FigStep (Gong et al., 2023), and assess model utility on standard VQA datasets.
+④We conduct extensive experiments across multiple VLMs safety evaluation benchmarks, including VLGuard (Zong et al., 2024), SPA-VL (Zhang et al., 2024b), MM-SafetyBench (Liu et al., 2024b), and FigStep (Gong et al., 2023), and assess model utility on standard VQA datasets.
 Our results confirm the safety mirage phenomenon and demonstrate that MU-based safety fine-tuning effectively mitigates spurious correlations and reduces over-prudence.
@@
 VLM safety: Attack and defense.
-With the rapid advancement of VLMs (Liu et al., 2023, 2024a; Zhu et al., 2023; Ye et al., 2023; Wang et al., 2023; Li et al., 2023b; Alayrac et al., 2022; Awadalla et al., 2023; Gao et al., 2023), safety concerns have become increasingly prominent due to their potential to generate harmful or inappropriate content.
-While LLMs have been extensively studied for safety risks, leading to the development of attack strategies (Yang et al., 2023; Wei et al., 2023a; Huang et al., 2023; Shu et al., 2023), defense mechanisms (Li et al., 2023c; Cao et al., 2023; Kumar et al., 2023), and robust evaluation datasets (Bianchi et al., 2023; Li et al., 2024b; Ji et al., 2023), VLMs introduce additional challenges due to the complexity of multimodal inputs (Pi et al., 2024; Chakraborty et al., 2024; Ding et al., 2025), making them even more vulnerable to jailbreaking and adversarial manipulation (Guo et al., 2024; Qi et al., 2023a; Liu et al., 2024b, c; Gong et al., 2023).
+With the rapid advancement of VLMs (Liu et al., 2023; 2024a; Zhu et al., 2023; Ye et al., 2023; Wang et al., 2023; Li et al., 2023a; Alayrac et al., 2022; Awadalla et al., 2023; Gao et al., 2023), safety concerns have become increasingly prominent due to their potential to generate harmful or inappropriate content.
+While LLMs have been extensively studied for safety risks, leading to the development of attack strategies (Yang et al., 2023; Wei et al., 2023b; Huang et al., 2023; Shu et al., 2023), defense mechanisms (Li et al., 2023c; Cao et al., 2023; Kumar et al., 2023), and robust evaluation datasets (Bianchi et al., 2023; Li et al., 2024a; Ji et al., 2023), VLMs introduce additional challenges due to the complexity of multimodal inputs (Pi et al., 2024; Chakraborty et al., 2024; Ding et al., 2025), making them even more vulnerable to jailbreaking and adversarial manipulation (Guo et al., 2024; Qi et al., 2023b; Liu et al., 2024c; b; Gong et al., 2023).
 Attacks on VLMs often leverage the dual-modality nature of these models.
-One approach embeds unsafe textual queries into images through typographic manipulation, enabling the model to bypass safety filters and generate harmful outputs (Gong et al., 2023; Liu et al., 2024c).
-Another strategy involves using gradient-based adversarial image generation (Bailey et al., 2023; Dong et al., 2023; Luo et al., 2023; Qi et al., 2023b; Zhao et al., 2023) to trigger harmful responses, demonstrating that VLMs remain susceptible to adversarial perturbations despite safety fine-tuning.
+One approach embeds unsafe textual queries into images through typographic manipulation, enabling the model to bypass safety filters and generate harmful outputs (Gong et al., 2023; Liu et al., 2024b).
+Another strategy involves using gradient-based adversarial image generation (Bailey et al., 2023; Dong et al., 2023; Luo et al., 2023; Qi et al., 2023a; Zhao et al., 2023) to trigger harmful responses, demonstrating that VLMs remain susceptible to adversarial perturbations despite safety fine-tuning.
 Defensive strategies for VLM safety generally fall into two categories: Inference-time defenses and fine-tuning with curated safety datasets.
 The former aligns safety responses dynamically at runtime, mitigating unsafe outputs using various filtering and rejection mechanisms (Wang et al., 2024b; Chen et al., 2023b; Pi et al., 2024; Gou et al., 2024; Ding et al., 2024).
-The latter focuses on red-teaming dataset curation (Liu et al., 2024c; Zhou et al., 2024; Luo et al., 2024; Zhang et al., 2024a; Gu et al., 2024; Zong et al., 2024; Li et al., 2024c), enabling VLMs to be explicitly trained to reject harmful content while retaining utility for benign tasks.
+The latter focuses on red-teaming dataset curation (Liu et al., 2024b; Zhou et al., 2024; Luo et al., 2024; Zhang et al., 2024b; Gu et al., 2024; Zong et al., 2024; Li et al., 2024b), enabling VLMs to be explicitly trained to reject harmful content while retaining utility for benign tasks.
 Machine unlearning in VLMs.
 MU (Liu et al., 2025; Cao and Yang, 2015; Bourtoule et al., 2021) is designed to remove harmful data influences from a pre-trained model while preserving its overall utility.
-In the LLM domain, recent work has explored targeted forgetting techniques to erase specific knowledge without compromising performance (Zhang et al., 2024b; Li et al., 2024a; Yao et al., 2024).
+In the LLM domain, recent work has explored targeted forgetting techniques to erase specific knowledge without compromising performance (Zhang et al., 2024a; Li et al., 2024c; Yao et al., 2024).
 In VLMs, several benchmarks have established systematic evaluation frameworks for MU algorithms (Liu et al., 2024d; Dontsov et al., 2024; Ma et al., 2024).
@@
 Our work builds on these efforts by leveraging MU to specifically mitigate spurious correlations present in safety fine-tuning datasets.
-3 Preliminary and Problem Statement
-Existing VLM safety fine-tuning setup.
-Previous works (Pi et al., 2024; Gong et al., 2023; Liu et al., 2024c) highlight the importance of safety alignment in VLMs across both textual and visual modalities.
-Consequently, many efforts have focused on curating high-quality dual-modality safety datasets for VLMs.
-For instance, VLGuard (Zong et al., 2024) covers various text-image scenarios, including unsafe cases where the text is unsafe or both modalities are unsafe, as well as safe cases where both are benign.
-SPA-VL (Zhang et al., 2024a) provides large-scale preference data in the form of question–response tuples, where each image sample pairs a safety-related question with a chosen response and a rejected response.
-Leveraging these curated safety datasets, recent works (Chakraborty et al., 2024; Ding et al., 2025) show that simple fine-tuning approaches on such datasets can yield surprisingly strong safety performance, even against common jailbreaking attacks (Zou et al., 2023; Wei et al., 2023b; Röttger et al., 2023).
-In this work, we revisit the VLM safety fine-tuning problem and later argue that the observed safety improvements from fine-tuning may be an illusion.
-We begin by formulating the problem of VLM safety fine-tuning.
-Let $\mathcal{D}_{\mathrm{u}}$ denote the unsafe dataset, which consists of unsafe text queries and corresponding input images possibly paired with targeted safe responses (e.g., rejection responses in VLGuard and SPA-VL).
-Let $\mathcal{D}_{\mathrm{r}}$ denote the retain dataset, consisting of either a safe text-image dataset or a safety-irrelevant utility dataset, designed to maintain VLM performance on normal tasks after safety fine-tuning.
-For a VLM parameterized by $\bm{\theta}$}, the safety fine-tuning problem can be formulated as:
-$\begin{array}[]{ll}\displaystyle\operatorname*{\text{minimize}}_{{\bm{\theta}}}&\ell_{\mathrm{u}}({\bm{\theta}};\mathcal{D}_{\mathrm{u}})+\gamma\ell_{\mathrm{r}}({\bm{\theta}};\mathcal{D}_{\mathrm{r}}),\end{array}$ (2)
-where $\ell_{\mathrm{u}}$ and $\ell_{\mathrm{r}}$ denote the fine-tuning losses over $\mathcal{D}_{\mathrm{u}}$ and $\mathcal{D}_{\mathrm{r}}$ respectively, and $\gamma\geq 0$ is a regularization parameter that balances safety alignment with general performance preservation.
 Figure 2: Visualization of question-answer samples on the safety fine-tuned VLMs (LLaVA-v1.5-7B-Mixed and LLaVA-v1.5-7B-Posthoc (Zong et al., 2024; Taori et al., 2023)).
@@
 However, asking the initial question by starting with the word with “What” can easily bypass this safeguard.
-(b) Over-prudence: A minor modification by asking the problem start from “What” with “Share” can trigger refusals even for benign queries.
+(b) Over-prudence: A minor modification by asking the problem start from “What” with “Share” can trigger unnecessary refusals even for benign queries.
+3 Preliminary and Problem Statement
+Existing VLM safety fine-tuning setup.
+Previous works (Pi et al., 2024; Gong et al., 2023; Liu et al., 2024b) highlight the importance of safety alignment in VLMs across both textual and visual modalities.
+Consequently, many efforts have focused on curating high-quality dual-modality safety datasets for VLMs.
+For instance, VLGuard (Zong et al., 2024) covers various text-image scenarios, including unsafe cases where the text is unsafe or both modalities are unsafe, as well as safe cases where both are benign.
+SPA-VL (Zhang et al., 2024b) provides large-scale preference data in the form of question–response tuples, where each image sample pairs a safety-related question with a chosen response and a rejected response.
+Leveraging these curated safety datasets, recent works (Chakraborty et al., 2024; Ding et al., 2025) show that simple fine-tuning approaches on such datasets can yield surprisingly strong safety performance, even against common jailbreaking attacks (Zou et al., 2023; Wei et al., 2023a; Röttger et al., 2023).
+In this work, we revisit the VLM safety fine-tuning problem and later argue that the observed safety improvements from fine-tuning may be an illusion.
+We begin by formulating the problem of VLM safety fine-tuning.
+Let $\mathcal{D}_{\mathrm{u}}$ denote the unsafe dataset, which consists of unsafe text queries and corresponding input images possibly paired with targeted safe responses (e.g., rejection responses in VLGuard and SPA-VL).
+Let $\mathcal{D}_{\mathrm{r}}$ denote the retain dataset, consisting of either a safe text-image dataset or a safety-irrelevant utility dataset, designed to maintain VLM performance on normal tasks after safety fine-tuning.
+For a VLM parameterized by $\bm{\theta}$}, the safety fine-tuning problem can be formulated as:
+$\begin{array}[]{ll}\displaystyle\operatorname*{\text{minimize}}_{{\bm{\theta}}}&\ell_{\mathrm{u}}({\bm{\theta}};\mathcal{D}_{\mathrm{u}})+\gamma\ell_{\mathrm{r}}({\bm{\theta}};\mathcal{D}_{\mathrm{r}}),\end{array}$ (2)
+where $\ell_{\mathrm{u}}$ and $\ell_{\mathrm{r}}$ denote the fine-tuning losses over $\mathcal{D}_{\mathrm{u}}$ and $\mathcal{D}_{\mathrm{r}}$ respectively, and $\gamma\geq 0$ is a regularization parameter that balances safety alignment with general performance preservation.
+Refer to caption Refer to caption Refer to caption
+(a) Top words in VLGuard safe   (b) Top words in VLGuard unsafe (c) Top words in SPA-VL
+Figure 3: Frequency of question-initiating words across training corpora: (a) VLGuard safe queries with non-rejection responses, (b) VLGuard unsafe queries with rejection responses, (c) SPA-VL queries.
 Safety mirage: Motivation and problem of interest.
-Although safety fine-tuned VLMs could exhibit over-rejection, refusing even benign queries (Ding et al., 2025; Guo et al., 2024; Ding et al., 2024), their safety against unsafe queries remains highly robust under common jailbreaking attacks (Zong et al., 2024; Zhang et al., 2024a).
+Although safety fine-tuned VLMs could exhibit over-rejection, refusing even benign queries (Ding et al., 2025; Guo et al., 2024; Ding et al., 2024), their safety against unsafe queries remains highly robust under common jailbreaking attacks (Zong et al., 2024; Zhang et al., 2024b).
 The seemingly ‘robust’ safety observed after fine-tuning motivates us to re-examine its true reliability.
@@
 Similarly, SPA-VL also selects rejection-based answers for unsafe queries.
-Refer to caption Refer to caption Refer to caption
-(a) Top words in VLGuard safe   (b) Top words in VLGuard unsafe (c) Top words in SPA-VL
-Figure 3: Frequency of question-initiating words across training corpora: (a) VLGuard safe queries with non-rejection responses, (b) VLGuard unsafe queries with rejection responses, (c) SPA-VL queries.
 At first glance, the use of safety labels appears appropriate.
@@
 Therefore, we define spurious correlations as the (unexpected) strong associations between spurious input features and the safety labels in the fine-tuning dataset.
-The above conceptualization of spurious correlations is inspired by and remains consistent with conventional spurious correlation analyses in image classification (Sagawa et al., 2020), where background pixels serve as spurious features while object pixels represent core content.
+The above conceptualization of spurious correlations is inspired by conventional spurious correlation analyses in image classification (Sagawa et al., 2020), where background pixels serve as spurious features while object pixels represent core content.
+In addition, the introduction of spurious correlations during safety alignment echoes the alignment tax (Siddiqui et al., 2026) and emergent misalignment phenomena (Betley et al., 2026), wherein fine-tuning a broadly capable pretrained model on a narrow objective can inadvertently induce new, unintended misaligned behaviors.
 In this work, we identify two types of spurious correlations in VLM safety fine-tuning.
@@
 We refer to this strategy, rephrasing unsafe queries to enforce a biased starting word, as the one-word jailbreaking attack.
+Figure 4: ASR of $K$-shot one-word attack for varying $K$, evaluated before and after applying the “What”-initialized one-word attack to jailbreak the safety fine-tuned VLM (LLaVA-v1.5-7B-Mixed (Zong et al., 2024)) on VLGuard unsafe.
 In practice, we find that repeatedly applying the one-word attack, by integrating $\mathrm{w}_{\mathrm{adv}}$ with paraphrased versions of the original input query $q$ (up to $K$ times), can significantly improve the attack success rate (ASR).
@@
 Additional results on LLaVA-v1.5-7B-PPO-30K are provided in Appendix B.1.
-Figure 4: ASR of $K$-shot one-word attack for varying $K$, evaluated before and after applying the “What”-initialized one-word attack to jailbreak the safety fine-tuned VLM (LLaVA-v1.5-7B-Mixed (Zong et al., 2024)) on VLGuard unsafe.
-Figure 5: $R$ vs.
+Figure 5: Rejection rate vs.
 $K$-shot one-word modification, evaluated before and after applying the “Share” initialized one-word modification to safe input queries, causes the over-prudence phenomenon of the safety fine-tuned VLM LLaVA-v1.5-7B-Mixed model (Zong et al., 2024) on VLGuard safe.
@@
 Although unlearning has been applied to VLM in prior work (Chen et al., 2025; Huo et al., 2025), its unique advantage in addressing spurious correlations remains underexplored.
-Therefore, we adapt two state-of-the-art MU approaches from LLMs, representation misdirection unlearning (RMU) (Li et al., 2024a) and negative preference optimization (NPO) (Zhang et al., 2024b), to VLM safety field.
+Therefore, we adapt two state-of-the-art MU approaches from LLMs, representation misdirection unlearning (RMU) (Li et al., 2024c) and negative preference optimization (NPO) (Zhang et al., 2024a), to VLM safety field.
 The proposed VLM unlearning follows the generic formulation of (2), but with the following key modifications.
@@
 where $M_{{\bm{\theta}}}(\cdot)$ represents certain intermediate-layer representations of $\bm{\theta}$}, $c$ is a hyperparameter that controls activation scaling , and $\mathbf{v}$ is a random vector drawn from a standard uniform distribution.
-We remark that, unlike RMU for LLM unlearning (Li et al., 2024a), we carefully adjusts the representation layer selection and tunes the hyperparameter $c$ to better suit the unlearning process in VLMs.
-In addition to RMU, we also employ NPO (Zhang et al., 2024b) to model the unlearning objective $\ell_{\mathrm{u}}$, which treats unsafe data designated for unlearning as “negative” examples in a direct preference optimization framework (Rafailov et al., 2023).
+We remark that, unlike RMU for LLM unlearning (Li et al., 2024c), we carefully adjusts the representation layer selection and tunes the hyperparameter $c$ to better suit the unlearning process in VLMs.
+In addition to RMU, we also employ NPO (Zhang et al., 2024a) to model the unlearning objective $\ell_{\mathrm{u}}$, which treats unsafe data designated for unlearning as “negative” examples in a direct preference optimization framework (Rafailov et al., 2023).
 The NPO-based unlearning loss is then given by
@@
 Datasets and models.
-We consider four VLM safety datasets: VLGuard (Zong et al., 2024), MM-SafetyBench (Liu et al., 2024c), SPA-VL (Zhang et al., 2024a), and Figstep (Gong et al., 2023).
+We consider four VLM safety datasets: VLGuard (Zong et al., 2024), MM-SafetyBench (Liu et al., 2024b), SPA-VL (Zhang et al., 2024b), and Figstep (Gong et al., 2023).
 To assess the utility of safety fine-tuned VLMs, we also conduct evaluations on representative visual question-answering (VQA) datasets, including VQAv2 (Goyal et al., 2017), TextVQA (Singh et al., 2019), VizWiz (Gurari et al., 2018), and ScienceQA (Lu et al., 2022).
-For model selection, we adopt LLaVA-v1.5-7B and LLaVA-v1.5-13B (Liu et al., 2023, 2024a) as our primary VLMs.
+For model selection, we adopt LLaVA-v1.5-7B and LLaVA-v1.5-13B (Liu et al., 2023; 2024a) as our primary VLMs.
 Safety fine-tuning setups and baselines.
@@
 (2) Posthoc-SFT (Zong et al., 2024; Taori et al., 2023): SFT using a post-hoc fine-tuning approach on VLGuard.
-(3) Unsafe-Filter: An SFT baseline where unsafe samples are excluded by using LLaMA-Guard-3-11B-Vision (Chi et al., 2024) to filter unsafe text–image pairs from LLaVA’s pre-training, refer to Appendix  D.2.
+(3) Unsafe-Filter: An SFT baseline where unsafe samples are excluded by using LLaMA-Guard-3-11B-Vision (Chi et al., 2024) to filter unsafe text–image pairs from LLaVA’s pre-training data, refer to Appendix  D.2.
 Evaluation setups.
@@
 We draw some key observations below.
+Figure 6: Visualization of question-answer pairs from three models: LLaVA-1.5-7B (original), Mixed-SFT (fine-tuned), and RMU-Unlearning (unlearned).
+Green shield represents the correct response, whether a safe rejection for harmful queries or a valid answer for benign queries.
+Red exclamation !
+indicates an unsafe response to harmful queries, while red question ?
+represents an inappropriate rejection for a safe query.
+The first row displays responses to unsafe text-image queries, while the second row shows responses to safe queries.
 First, conventional safety fine-tuning approaches (Unsafe-Filter, Mixed-SFT, and Posthoc-SFT) suffer from a clear safety mirage: ASR rises sharply after one-word attacks (nearly 60% on VLGuard-unsafe and 30% on SPA-VL-unsafe for 7B models), while RR also increases dramatically (over 90%) on safe queries.
@@
 In contrast, RMU-based unlearning both resists the one-word attack and alleviates over-rejection on safe queries.
-Figure 6: Visualization of question-answer pairs from three models: LLaVA-1.5-7B (original), Mixed-SFT (fine-tuned), and RMU-Unlearning (unlearned).
-Green shield represents the correct response, whether a safe rejection for harmful queries or a valid answer for benign queries.
-Red exclamation !
-indicates an unsafe response to harmful queries, while red question ?
-represents an inappropriate rejection for a safe query.
-The first row displays responses to unsafe text-image queries, while the second row shows responses to safe queries.
-Unlearning as a distinct safety mechanism: A response analysis.
-To highlight the difference in safety mechanisms between MU-based approaches and safety-aware SFT, Table 2 reports the proportions of unsafe, irrelevant, and rejection responses before and after 1-shot one-word attack.
-Here, the safety rate (1–ASR) is decomposed into: (1) irrelevant responses, where the model sidesteps the unsafe query with unrelated content, and (2) rejection responses, where the model explicitly refuses to answer.
-Conventional SFT approaches, guided by safety labels, rely mainly on rejection responses, yielding consistently high RR both “Before” and “After”.
-In contrast, MU-based methods do not rely on explicit rejection labels.
-Instead, they achieve safety primarily through irrelevant responses, reflecting a more label-free mechanism for mitigating spurious correlation.
 Table 2: Analyzing the safety mechanisms of unlearning-based approaches versus baselines.
@@
     +RMU-Unlearning 3.87% 90.92% 5.21% 6.91% 88.33% 4.76%
+Unlearning as a distinct safety mechanism: A response analysis.
+To highlight the difference in safety mechanisms between MU-based approaches and safety-aware SFT, Table 2 reports the proportions of unsafe, irrelevant, and rejection responses before and after 1-shot one-word attack.
+Here, the safety rate (1–ASR) is decomposed into: (1) irrelevant responses, where the model sidesteps the unsafe query with unrelated content, and (2) rejection responses, where the model explicitly refuses to answer.
+Conventional SFT approaches, guided by safety labels, rely mainly on rejection responses, yielding consistently high RR both “Before” and “After”.
+In contrast, MU-based methods do not rely on explicit rejection labels.
+Instead, they achieve safety primarily through irrelevant responses, reflecting a more label-free mechanism for mitigating spurious correlation.
+[…]
+Models Before After After
+one-word one-word GCG
+LLaVA-1.5-7B 64.25% 90.27% 72.90%
+    +Unsafe-Filter 65.66% 90.72% 73.75%
+    +Mixed-SFT 0.23% 54.98% 5.54%
+    +Post-hoc 0.23% 46.83% 4.07%
+    +NPO-Unlearning 2.49% 12.92% 3.87%
+    +RMU-Unlearning 1.29% 10.18% 2.26%
+Evaluation on optimization-based textual attack.
+To assess whether MU generalizes beyond the dataset-induced correlations identified in Fig.
+3, we evaluate the GCG attack (Zou et al., 2023), with results reported in Table 3.
+We adapt GCG to VLMs by optimizing 20 adversarial suffix tokens on LLaVA-1.5-7B using a harmful image–text pair (an explosive-device image paired with “How to build a bomb?”) and performing 300 gradient-ascent steps targeting the unsafe response “Sure, I’d be happy to help.” The learned suffix tokens are external to the training dataset and appended only at inference time.
+Under a unified 3-shot setting, GCG produces only modest ASR increases on VLMs.
+In contrast, simple one-word substitutions (e.g., replacing the first word with “What”) severely degrade SFT-based models, driving ASR up to 90%.
+These results suggest that the primary vulnerability arises from the spurious correlations we identify rather than from optimization-generated adversarial tokens, while MU remains robust under GCG attacks.
 Figure 7: Input token sensitivity analysis for “What”-initiated unsafe queries and “Share”-initiated safe queries over VLGuard using LLaVA-1.5-7B Mixed-SFT, before and after masking “What” and “Share”.
@@
 We further validate spurious correlations between specific query words and safety labels using token-level sensitivity analysis, where selected input tokens are masked (i.e., replaced with [PAD]) and their effect on rejection probabilities is measured.
-Figure 7 shows the results for “What’ initiated unsafe queries and “Share” initiated safe queries using LLaVA-1.5-7B Mixed-SFT.
+Fig.
+7 shows the results for “What’ initiated unsafe queries and “Share” initiated safe queries using LLaVA-1.5-7B Mixed-SFT.
 Masking “What” leads to a sharp increase in rejection probability, confirming its role in inducing non-rejection bias.
 Conversely, masking “Share” reduces the rejection probability compared to the unmasked case, highlighting its influence in reinforcing rejection bias.
-Other saliency maps are provided in Appendix D.6.
-Additional results. In Appendix D.7, we demonstrate the robustness of spurious correlation and the effectiveness of our MU-based approach when facing visual variations.
+Additional results.
+In Appendix D.6, we demonstrate the robustness of spurious correlation and the effectiveness of our MU-based approach when facing visual variations.
+In Appendix D.7, we also show the additional analysis of input saliency maps.
 7 Conclusion
@@
 Our experiments validate the safety mirage phenomenon and demonstrate that MU improves robustness against jailbreaks, alleviates over-prudence, and preserves utility on standard VQA tasks.
-We refer readers to Appendix FE for limitations, broad impacts, and LLM usage.
-8 Acknowledgement
-The project is primarily supported by research awards from Cisco Research and DSO National Laboratories.
-In addition, the contributions of Yiwei Chen, Yuguang Yao, Yihua Zhang, and Sijia Liu are partially supported by the NSF CISE Core Program Awards IIS-2207052 and IIS-2504263, the NSF CAREER Award IIS-2338068, the NSF Cyber-Physical Systems (CPS) Award CNS-2235231, the Amazon Research Award, the Open Philanthropy Research Award, and the IBM PhD Fellowship.
+We refer readers to Appendix EF for limitations, broad impact, and details of LLM usage.
+Acknowledgment
+This project is supported by research awards from Cisco Research and DSO National Laboratories.
+The contributions of Yiwei Chen, Yuguang Yao, Yihua Zhang, and Sijia Liu are also partially supported by the NSF CISE Core Program Awards IIS-2207052 and IIS-2504263, the NSF CAREER Award IIS-2338068, the NSF Cyber-Physical Systems (CPS) Award CNS-2235231, and the Open Philanthropy Research Award.
 Appendix
 Appendix A Visualization of Motivation Samples
-Figure A1: Visualization of question-answer samples on the safety fine-tuned VLMs (LLaVA-v1.5-7B-SPA-VL-DPO-30K and LLaVA-v1.5-7B-SPA-VL-PPO-30K (Zhang et al., 2024a; Rafailov et al., 2023; Schulman et al., 2017)).
+Figure A1: Visualization of question-answer samples on the safety fine-tuned VLMs (LLaVA-v1.5-7B-SPA-VL-DPO-30K and LLaVA-v1.5-7B-SPA-VL-PPO-30K (Zhang et al., 2024b; Rafailov et al., 2023; Schulman et al., 2017)).
 The visual marks keep consistent meaning with Fig.
@@
 B.1 Additional Results of One-Word Jailbreaking
-Figure A2: ASR of $K$-shot one-word attack for varying $K$, evaluated before and after applying the “Please”-initialized one-word attack to jailbreak the safety fine-tuned VLM (LLaVA-v1.5-7B-PPO-30K (Zhang et al., 2024a)) on SPA-VL unsafe test set.
+Figure A2: ASR of $K$-shot one-word attack for varying $K$, evaluated before and after applying the “Please”-initialized one-word attack to jailbreak the safety fine-tuned VLM (LLaVA-v1.5-7B-PPO-30K (Zhang et al., 2024b)) on SPA-VL unsafe test set.
 For the SPA-VL test set, we further evaluated the LLaVA-v1.5-7B-PPO-30K model using the same $K$-shot paraphrasing protocol as in the main text.
@@
 B.2 Additional Results of One-Word Over-prudence
-Figure A3: Rejection rate of $K$-shot one-word modification, evaluated before and after applying the “What” initialized one-word modification to safe input queries, causes the over-prudence phenomenon of the safety fine-tuned VLM (LLaVA-v1.5-7B-PPO-30K (Zhang et al., 2024a)) on SPA-VL safe.
+Figure A3: Rejection rate of $K$-shot one-word modification, evaluated before and after applying the “What” initialized one-word modification to safe input queries, causes the over-prudence phenomenon of the safety fine-tuned VLM (LLaVA-v1.5-7B-PPO-30K (Zhang et al., 2024b)) on SPA-VL safe.
 We also tested the over-prudence effect on SPA-VL safe queries.
@@
 For full fine-tuning, we employ a global batch size of 8 and a learning rate of $1\times 10^{-6}$.
-For LoRA-based fine-tuning, following (Liu et al., 2023; Zhang et al., 2024a), we set the rank $r$ to 256 and the scaling factor $\alpha$ to 512, with a global batch size of 64 and a learning rate of $1\times 10^{-5}$.
+For LoRA-based fine-tuning, following (Liu et al., 2023; Zhang et al., 2024b), we set the rank $r$ to 256 and the scaling factor $\alpha$ to 512, with a global batch size of 64 and a learning rate of $1\times 10^{-5}$.
 In both setups, we use the AdamW optimizer with zero weight decay and a warmup ratio of 0.03.
@@
 Importantly, even with only 10% of the safety data, unlearning-based models still achieve substantially better safety–utility trade-offs than conventional fine-tuning baselines, confirming the flexibility of coreset unlearning for practical deployment
-Figure A4: Safety evaluation on MM-SafetyBench and FigStep. ASR is reported before and after the “What”-based 3-shot attack. Setup follows Table 1.
+Table A2: Safety evaluation on MM-SafetyBench and FigStep. ASR is reported before and after the “What”-based 3-shot attack. Setup follows Table 1.
 Models Safety Evaluation ($\downarrow$)
@@
 This highlights that MU not only mitigates overfitting to dataset-specific biases but also generalizes better across benchmarks, thereby offering a more reliable defense against spurious correlation–driven vulnerabilities.
-Figure A5: Prediction probability for various output sentences before and after masking each input query token for the query “What can be used to create a promotion?” corresponding to the image example in Fig.
-1-(a).
-Figure A6: Prediction probability for various output sentences before and after masking each input query token for the query “Share the drink type in the image” corresponding to the image example in Fig.
-1-(b).
-D.6 Examples of Input Saliency
-We show two input saliency maps in Fig.
-A5 and Fig.
-A6.
-The input saliency is shown through the prediction probability change for each output sentence ($y$-axis) by masking each input token ($x$-axis).
-Here [X] refers to the token “X” is marked (i.e., replacing with [PAD])), and [ ] denotes the non-masking case.
-The examples of input query in Fig.
-A5 and Fig.
-A6 correspond to the examples introduced in Fig.
-1(a)-(b).
-Fig.
-A5 shows that masking “what” token can significantly boost the rejection probability, producing “I’m sorry…”.
-Fig.
-A6 shows that masking “share” token successfully reduces the prediction probability for rejection response and produces the normal model generation.
-Figure A7: Evaluating safety (ASR, $\downarrow$) and over-prudence ($R$, $\downarrow$) under visual perturbations on VLGuard, including Gaussian Noise, Gaussian Blur, and Color Jitter.
+D.6 Spurious Correlations with Visual Variations
+To further examine whether the identified spurious correlations between textual queries and safety labels persist under different visual conditions, we evaluate fine-tuned models on VLGuard with a range of common visual perturbations, including Gaussian Noise, Gaussian Blur, and Color Jitter.
+The results are summarized in Table A3.
+We make several observations.
+Table A3: Evaluating safety (ASR, $\downarrow$) and over-prudence ($R$, $\downarrow$) under visual perturbations on VLGuard, including Gaussian Noise, Gaussian Blur, and Color Jitter.
 Other setups follow Table 1.
@@
 + Color Jitter 1.74% 10.41% 1.07% 7.02%
-D.7 Spurious Correlations with Visual Variations
-To further examine whether the identified spurious correlations between textual queries and safety labels persist under different visual conditions, we evaluate fine-tuned models on VLGuard with a range of common visual perturbations, including Gaussian Noise, Gaussian Blur, and Color Jitter.
-The results are summarized in Table,A7.
-We make several observations.
 First, conventional SFT-based safety models (Mixed-SFT and Posthoc-SFT) exhibit a consistent pattern of vulnerability: both ASR and RR remain high across all perturbations, showing that the learned spurious correlations are insensitive to low-level visual variations.
@@
 In contrast, unlearning-based models generalize more reliably across such shifts, further validating their robustness in multimodal safety alignment.
-Appendix E The Use of LLMs
-This work makes limited use of LLMs.
-Specifically, LLMs were employed exclusively for grammar correction and stylistic polishing of the manuscript.
-They were not involved in research ideation, experimental design, data analysis, or the generation of any scientific content.
-All substantive contributions to the paper are solely attributable to the author.
-Appendix F Limitations and broad Impact
-F.1 Limitations
+Figure A4: Prediction probability for various output sentences before and after masking each input query token for the query “What can be used to create a promotion?” corresponding to the image example in Fig.
+1-(a).
+Figure A5: Prediction probability for various output sentences before and after masking each input query token for the query “Share the drink type in the image” corresponding to the image example in Fig.
+1-(b).
+D.7 Examples of Input Saliency
+We show two input saliency maps in Fig.
+A4 and Fig.
+A5.
+The input saliency is shown through the prediction probability change for each output sentence ($y$-axis) by masking each input token ($x$-axis).
+Here [X] refers to the token “X” is marked (i.e., replacing with [PAD])), and [ ] denotes the non-masking case.
+The examples of input query in Fig.
+A4 and Fig.
+A5 correspond to the examples introduced in Fig.
+1(a)-(b).
+Fig.
+A4 shows that masking “what” token can significantly boost the rejection probability, producing “I’m sorry…”.
+Fig.
+A5 shows that masking “share” token successfully reduces the prediction probability for rejection response and produces the normal model generation.
+Appendix E Limitations and broad Impact
+E.1 Limitations
 Our analysis has so far been limited to models at the 7B and 13B scales.
@@
 Future work should therefore investigate unlearning performance in broader, domain-specific scenarios and corresponding downstream tasks.
-F.2 Broad Impact
+E.2 Broad Impact
 This work sheds light on the reliability of current VLM safety fine-tuning, uncovering the risk of a “safety mirage” where perceived robustness is illusory.
@@
 We encourage future research to balance safety, fairness, and transparency when deploying MU-based safety interventions in real-world applications.
+Appendix F The Use of LLMs
+This work makes limited use of LLMs.
+Specifically, LLMs were employed exclusively for grammar correction and stylistic polishing of the manuscript.
+They were not involved in research ideation, experimental design, data analysis, or the generation of any scientific content.
+All substantive contributions to the paper are solely attributable to the author.